### v1

# Adam Exploits $\ell_{\infty}$-geometry of Loss Landscape via Coordinate-wise Adaptivity

Abstract
Adam outperforms SGD when training language models. Yet this advantage is not well-understood theoretically – previous convergence analysis for Adam and SGD mainly focuses on the number of steps $T$ and is already minimax-optimal in non-convex cases, which are both $\widetilde{O}(T^{-1/4})$. In this work, we argue that the exploitation of nice $\ell_{\infty}$-geometry is the key advantage of Adam over SGD. More specifically, we give a new convergence analysis for Adam under novel assumptions that loss is smooth under $\ell_{\infty}$-geometry rather than the more common $\ell_{2}$-geometry, which yields a much better empirical smoothness constant for GPT-2 and ResNet models. Our experiments confirm that Adam performs much worse when the favorable $\ell_{\infty}$-geometry is changed while SGD provably remains unaffected. We also extend the convergence analysis to blockwise Adam under novel blockwise smoothness assumptions.
1 Introduction
Large language models (LLMs) have gained phenomenal capabilities as their scale grows (Radford et al., 2019; Kaplan et al., 2020; Brown et al., 2020; Zhang et al., 2022a; Touvron et al., 2023; OpenAI, 2023; Reid et al., 2024). However, pre-training LLMs are incredibly time-consuming. Adaptive Momentum Estimation ($\mathtt{Adam}$)(Kingma & Ba, 2014) is the current to-go optimization algorithm for LLMs due to its fast convergence. In contrast, $\mathtt{SGD}$, a popular and arguably the simplest optimizer, optimizes language model loss much slower than $\mathtt{Adam}$.
However, the optimization benefit of $\mathtt{Adam}$ over $\mathtt{SGD}$ cannot be explained by existing theory. Existing convergence analyses for $\mathtt{Adam}$ and $\mathtt{SGD}$ focus on the dependence on the number of steps under assumptions on the smoothness and gradient bounds of the loss (Défossez et al., 2022), and it has been shown that both $\mathtt{Adam}$ and $\mathtt{SGD}$ achieve the minimax convergence rate $\widetilde{O}(T^{-1/4})$ in the non-convex settings (Arjevani et al., 2023). Thus according to the theory, in the worst case, $\mathtt{SGD}$ would be more desirable compared to $\mathtt{Adam}$ because they have the same convergence rate, and yet $\mathtt{Adam}$ is less memory-efficient due to its coordinate-wise adaptivity, which needs to store the empirical moving average of second-order moments of past stochastic gradients. Therefore, we hypothesize that the coordinate-wise adaptivity in $\mathtt{Adam}$ is exploiting some unknown properties of LLMs which $\mathtt{SGD}$ cannot make use of.
Towards this end, we identified a significant difference between $\mathtt{Adam}$ and $\mathtt{SGD}$ in this paper. This difference, often ignored in previous works, is that $\mathtt{SGD}$ is rotation-invariant, while $\mathtt{Adam}$ is only permutation-invariant (see definitions in Section 2). Intuitively, this means if we rotate the loss landscape, the optimization trajectory of $\mathtt{SGD}$ would be the same (up to some rotation), while the trajectory of $\mathtt{Adam}$ could be completely different. If $\mathtt{Adam}$ optimizes much slower more slowly after rotation, then it suggests $\mathtt{Adam}$ is exploiting some non-rotational-invariant properties of the loss function, which is not captured by standard smoothness assumptions in the convergence analysis.
Figure 1 summarizes our findings by comparing $\mathtt{Adam}$ on the original and rotated loss. The performance of $\mathtt{Adam}$ on rotated loss does become much worse than $\mathtt{Adam}$ on the original loss. We also test a memory-efficient and rotational-invariant variant of $\mathtt{SGD}$, $\mathtt{AdaSGD}$ (Wang & Wiens, 2020) (defined in Algorithm 2) [^1] There is one small difference. We use an exponential average of the gradient for ${\bm{m}}_{t}$ instead of momentum. Our definition makes $\mathtt{AdaSGD}$ the same as $\mathtt{Adam}$ in a one-dimensional problem.. Surprisingly, the rotated $\mathtt{Adam}$ performs even much worse than the $\mathtt{SGD}$ variant. The results suggest it is impossible to explain the superior optimization performance of Adam over SGD just using rotationally invariant assumptions on the loss function, which raises the natural question,
 What non-rotation-invariant properties of loss functions enable $\mathtt{Adam}$ to converge faster than $\mathtt{SGD}$?
We hypothesize that the $\ell_{2}$ lipschitzness of loss gradient does not provide a tight-enough characterization of loss landscape of deep learning models in practice, such that we can separate $\mathtt{Adam}$ and other rotational-invariant algorithms. Inspired by the similarity between $\mathtt{Adam}$ and $\mathtt{SignGD}$ and the fact that $\mathtt{SignGD}$ is the normalized steepest descent with respect to $\ell_{\infty}$ norm, we propose to use $\ell_{\infty}$-norm related smoothness as a better tool to analyze $\mathtt{Adam}$. In particular, our main results use the $(1,1)$-norm of Hessian of loss divided by variable dimension $d$ in replacement of the spectral norm of Hessian as the smoothness measure, and prove a convergence rate of $O(\frac{1}{\sqrt{T}})$ for $\mathtt{Adam}$ without noise, or $O((\frac{\log T}{T})^{1/4})$ with noise. Our results have the same dependence on $T$ as previous results, but much smaller smoothness constant when we measure it empirically. We empirically verify that $(1,1)$-norm of Hessian positively correlates with final training loss of $\mathtt{Adam}$ on both synthetic tasks like quadratic loss and real tasks like training GPT2 on OpenWebText and ResNet on CIFAR10.
We summarize our contributions below:
  1. 1.
    We show by experiments that the empirical optimization advantage of $\mathtt{Adam}$ over $\mathtt{SGD}$ can not explained solely under rotation-invariant assumptions. (Figure 1)
  2. 2.
    We propose a new complexity metric for the optimization problem, which is the $(1,1)$-norm of the Hessian matrix of loss, $\left\|\nabla^{2}L(x)\right\|_{1,1}$. We present a novel convergence result for $\mathtt{Adam}$ depending on this metric in the case of $\beta_{1}=0$. (Theorem 3.5 )
  3. 3.
    We further generalize the theoretical analysis for $\mathtt{Adam}$ to blockwise $\mathtt{Adam}$ (Algorithm 3) whose convergence rate can be characterized by a novel smoothness measure (Theorem 3.11). $\mathtt{Adam}$ and $\mathtt{AdaSGD}$ are two notable examples of blockwise $\mathtt{Adam}$. In $\mathtt{Adam}$, all blocks are of size $1$. In $\mathtt{AdaSGD}$, there is only one block.
  4. 4.
    We empirically verify that when $\mathtt{Adam}$ converges more slowly on the rotated loss, the $(1,1)$-norm of Hessian also increases, which suggests that our new complexity metric for $\mathtt{Adam}$’s convergence is practically relevant. (Section 4). [^2] The code is available at https://github.com/mohamad-amin/adam-coordinate-adaptivity.
Figure 1: Training and evaluation losses of $\mathtt{Adam}$, $\mathtt{AdaSGD}$ and $\mathtt{SGD}$ on GPT-2. $\mathtt{rotated\ Adam}$ means running $\mathtt{Adam}$ on a rotated loss. $\mathtt{Adam}$ on the original loss converges the fastest as expected. But convergence of $\mathtt{Adam}$ on a rotated loss is much slower, notably even worse than $\mathtt{AdaSGD}$.
Algorithm 1 Adam
$\beta_{1},\beta_{2},\epsilon\geq 0$, total steps $T$, learning rate $\{\eta_{t}\}_{t=1}^{T}$, $\epsilon$, initial ${\bm{m}}_{0},v_{0}$
initialization ${\bm{x}}_{0}$, stochastic loss functions $\{L_{t}\}_{t=1}^{T}$
$v_{0,i}\leftarrow v_{0}$
for $t=1,2,\cdots,T$ :
   $g_{t,i}\leftarrow\nabla_{i}L_{t}({\bm{x}}_{t-1})$
   $m_{t,i}\leftarrow\beta_{1}m_{t-1,i}+(1-\beta_{1})g_{t,i}$
   $v_{t,i}\leftarrow\beta_{2}v_{t-1,i}+(1-\beta_{2})g_{t,i}^{2}$
   $x_{t,i}\leftarrow x_{t-1,i}-\eta_{t}\frac{m_{t,i}}{\sqrt{v_{t,i}+\epsilon}}$
return ${\bm{x}}_{T}$
Algorithm 2 AdaSGD
$\beta_{1},\beta_{2},\epsilon\geq 0$, total steps $T$, learning rate $\{\eta_{t}\}_{t=1}^{T}$, initial ${\bm{m}}_{0},v_{0}$
initialization ${\bm{x}}_{0}$, stochastic loss functions $\{L_{t}\}_{t=1}^{T}$
for $t=1,2,\cdots,T$ :
   $g_{t,i}\leftarrow\nabla_{i}L_{t}({\bm{x}}_{t-1})$
   $m_{t,i}\leftarrow\beta_{1}m_{t-1,i}+(1-\beta_{1})g_{t,i}$
   $v_{t}\leftarrow\beta_{2}v_{t-1}+(1-\beta_{2})(\left\|{\bm{g}}_{t}\right\|_{2}^{2}/d)$
   $x_{t,i}\leftarrow x_{t-1,i}-\eta_{t}\frac{m_{t,i}}{\sqrt{v_{t}+\epsilon}}$
return ${\bm{x}}_{T}$
2 Preliminaries
Notations. For ${\bm{x}}\in\mathbb{R}^{d}$, we define the vector $p$-norm $\|{\bm{x}}\|_{p}$ as $(\sum_{i=1^{d}}x_{i}^{p})^{1/p}$ for $p\in[1,\infty]$. For a matrix $\bm{A}\in\mathbb{R}^{d_{1}\times d_{2}}$, its $(1,1)$-norm $\|\bm{A}\|_{1,1}$ is defined as $\sum_{i=1}^{d_{1}}\sum_{j=1}^{d_{2}}\left|A_{i,j}\right|$ and its operator norm induced by vector $p$-norm $\|\cdot\|_{p}$ as $\sup_{{\bm{x}}\in\mathbb{R}^{d}}\frac{\|\bm{A}{\bm{x}}\|_{q}}{\|{\bm{x}}\|_{p}}$, denoted by $\|\bm{A}\|_{p}$, where $\frac{1}{q}+\frac{1}{p}=1$ and $\|\cdot\|_{q}$ is the dual norm of $\|\cdot\|_{p}$. For a deterministic loss function $L({\bm{x}})$, we consider optimization over $L$ with only access independent stochastic functions $\{L_{t}({\bm{x}})\}_{t=1}^{T}$ such that $\mathbb{E}L_{t}({\bm{x}})=L({\bm{x}})$ for any input ${\bm{x}}\in\mathbb{R}^{d}$.
Rotation. For an invertible function $\mathcal{T}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, $\mathcal{T}$ is a rotating transformation if there exists an orthogonal matrix $\bm{T}\in\mathbb{R}^{d\times d}$ such that $\mathcal{T}({\bm{x}})=\bm{T}{\bm{x}}$. $\mathcal{T}$ is a permutating transformation if there exists a permutation $\pi:[d]\rightarrow[d]$ such that $\mathcal{T}({\bm{x}})=[x_{\pi(1)},\dots,x_{\pi(d)}]^{\top}$. A permutating transformation is always a rotating transformation. We will use $\mathcal{R}$ to denote a rotating transformation.
Definition 2.1.
For initialization ${\bm{x}}_{0}$ and stochastic losses $\{L_{t}\}_{t=1}^{T}$, we can get ${\bm{x}}_{t}$ when running algorithm $A$ on $({\bm{x}}_{0},\{L_{t}\}_{t=1}^{T})$. For a transformation $\mathcal{T}$, we can also get $\tilde{{\bm{x}}}_{t}$ when running $A$ with the same hyperparameters on $(\tilde{{\bm{x}}}_{0},\{\tilde{L}_{t}\}_{t=1}^{T})$ with $\tilde{{\bm{x}}}_{0}=\mathcal{T}^{-1}({\bm{x}}_{0})$ and $\tilde{L}_{t}=L_{t}\circ\mathcal{T}$.
An algorithm $A$ is invariant w.r.t. $\mathcal{T}$ if it always holds that $\tilde{{\bm{x}}}_{t}=\mathcal{T}^{-1}({\bm{x}}_{t})$ for any hyperparameters, initialization and stochastic losses. An algorithm $A$ is rotation invariant if it is invariant w.r.t. any rotating transformation $\mathcal{R}$. And $A$ is permutation invariant if it is invariant w.r.t. any permutating transformation.
The following Theorem 2.2 shows the difference between $\mathtt{Adam}$ and $\mathtt{AdaSGD}$, whose proof is in Appendix B.
Theorem 2.2.
$\mathtt{SGD}$ and $\mathtt{AdaSGD}$ are rotation-invariant. $\mathtt{Adam}$ and $\mathtt{SignGD}$ are only permutation-invariant.
3 Main Results: Convergence Rates of $\mathtt{Adam}$
In this section, we present our main theoretical results, starting with a convergence analysis of $\mathtt{Adam}$ for stochastic smooth loss with coordinate-wise gradient noise (Theorem 3.5). We allow non-convex losses and thus the convergence is measured by the $\ell_{1}$ norm of the gradient. For deterministic loss, our best convergence rate (Theorem 3.2) is achieved by $\mathtt{SignGD}$ ($\mathtt{Adam}$ with $\beta_{1}=\beta_{2}=0$). For stochastic loss with bounded gradient noise variance, our best rate (Corollary 3.6) is achieved by $\mathtt{RMSProp}$ ($\mathtt{Adam}$ with $\beta_{1}=0$ and $\beta_{2}\in[0,1]$).
Then we extend our analysis of $\mathtt{Adam}$ to more general blockwise $\mathtt{Adam}$ (Theorem 3.11), which contains both $\mathtt{Adam}$ and $\mathtt{AdaSGD}$ as special cases. We also comes up with novel smoothness measures (Definition 3.9) corresponding to the set of blocks used in blockwise $\mathtt{Adam}$.
Similar to previous works (Défossez et al., 2022), our analysis could be extended to the most general case of $\mathtt{Adam}$, where both $\beta_{1}$, $\beta_{2}$ are non-zero, but the rate becomes strictly worse than the $\mathtt{RMSProp}$ (the case of $\beta_{1}=0$), as there will be some extra polynomials of $\frac{1}{1-\beta_{1}}$. We decide not to include result for the most general case, on one hand for ease of presentation, and on the other hand, because such result could explain the optimization benefit of momentum ($\beta_{1}>0$) in practice and does not add any insight on the benefit of $\mathtt{Adam}$. We hypothesis that we are missing some important features of loss landscape of transformers in the theoretical assumptions and we leave this for future work.
3.1 Warmup: $\mathtt{SignGD}$ ($\beta_{1}=\beta_{2}=0$)
In this section, we use the convergence analysis for $\mathtt{SignGD}$ ($\mathtt{Adam}$ with $\beta_{1}=\beta_{2}=0$) as a warm-up and illustrate how $\mathtt{Adam}$ could benefit from a non-rotational invariant property of the loss landscape, which in particular is the $\ell_{\infty}$ smoothness. The key observation here is that $\mathtt{SignGD}$ is the normalized steepest descent with respect to $\ell_{\infty}$ norm (see (Xie & Li, 2024)), and thus it is more natural to analyze its convergence using $\ell_{\infty}$-norm-related geometry of the loss.
Definition 3.1.
[…]
Theorem 3.2.
Let $L$ be $H$-smooth with respect to $\ell_{\infty}$ norm and $\{{\bm{x}}_{t}\}_{t=1}^{T}$ be the iterates of $\mathtt{SignGD}$ ($\mathtt{Adam}$ with $\beta_{1}=\beta_{2}=0$) on $L$ with initialization ${\bm{x}}_{0}$ and learning rate $\eta$, it holds that
$\min_{1\leq t\leq T}\left\|\nabla L({\bm{x}}_{t})\right\|_{1}\leq\frac{L({\bm{x}}_{0})-\min L}{T\eta}+\frac{H\eta}{2}$
if we choose $\eta=\sqrt{\frac{2(L({\bm{x}}_{0})-\min L)}{TH}}$, then $\min_{1\leq t\leq T}\left\|\nabla L({\bm{x}}_{t})\right\|_{1}\leq\sqrt{\frac{2H(L({\bm{x}}_{0})-\min L)}{T}}$.
3.2 Main result: $\mathtt{RMSProp}$ ($\beta_{1}=0,\beta_{2}\in[0,1]$)
It is well-known that $\mathtt{SignGD}$ might not converge in the stochastic case as the expectation of descent direction for mini-batch loss may not be a descent direction, and $\mathtt{RMSProp}$ is proposed to address this issue by using a moving average of the squared gradient per coordinate to reduce the coorleation between the denominator and the numerator, thus making the expected update direction less biased (Hinton et al., 2012). In this subsection we formalize the above intuition and show indeed a positive $\beta_{2}$ in $\mathtt{Adam}$ helps convergence in the stochastic case. The main challenges here are from both lower bounding the first-order term and upper bounding the second-order term in the modified descent lemma (the counterpart of Appendix A for $\mathtt{RMSProp}$).
$L({\bm{x}}_{t})-L({\bm{x}}_{t-1})\leq-\eta_{t}\nabla L({\bm{x}}_{t})^{\top}\frac{{\bm{g}}_{t}}{\sqrt{{\bm{v}}_{t}+\epsilon}}+\frac{H}{2}\eta_{t}^{2}\left\|\frac{{\bm{g}}_{t}}{\sqrt{{\bm{v}}_{t}+\epsilon}}\right\|_{\infty}^{2}$
We can only upper bound $\left\|\frac{{\bm{g}}_{t}}{\sqrt{{\bm{v}}_{t}+\epsilon}}\right\|_{\infty}^{2}$ by $\frac{1}{1-\beta_{2}}$ without more fine-grained analysis on the relationship between gradients in each step, which will greatly hurt the dependence of convergence rate on $1-\beta_{2}$. However, even though the update at step $t$ for one specific coordinate $i$ can be as large as $\frac{1}{\sqrt{1-\beta_{2}}}$ with some very large $g_{t,i}$, the average moving speed for each coordinate should be close to $1$. Therefore, we introduce a slightly stronger definition in Definition 3.3, which allows us to decompose the second order term into each coordinate according to Lemma 3.12. It also facilitates the analysis for the coordinate-wise first order term. We note this definition also appears in Assumption 2.3 of the concurrent work Maladkar et al. (2024).
Definition 3.3.
For any $\mathbf{H}=(H_{1},\ldots,H_{d})\in\mathbb{R}^{d}$, we say a function $L$ is $\mathbf{H}$-smooth coordinate-wisely w.r.t. $\ell_{\infty}$ norm , iff for any $i\in[d]$, ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d}$, $\left|\nabla_{i}L({\bm{x}})-\nabla_{i}L({\bm{y}})\right|\leq H_{i}\left\|{\bm{x}}-{\bm{y}}\right\|_{\infty}.$
By definition, $\mathbf{H}$-smoothness coordinate-wisely w.r.t. $\ell_{\infty}$ norm implies $\sum_{i=1}^{d}H_{i}$ smoothness w.r.t. $\ell_{\infty}$ norm. We also need 3.4 to measure the influence of noise in stochastic setting.
Assumption 3.4 (Coordinate-wise noise).
There exist constants $\sigma_{i}$ such that
$\mathbb{E}\left(\nabla_{i}L_{t}({\bm{x}})-\nabla_{i}L({\bm{x}})\right)^{2}\leq\sigma_{i}^{2}$
for any $i\in[d]$, $t\in\mathbb{N}$ and ${\bm{x}}\in\mathbb{R}^{d}$.
Due to the limitation of space, we only present the main result here. The sketch of the proof is presented in Section 3.4. We present the complete proof for the generalized blockwise Adam algorithm in Appendix C. The proof incorporates some key steps from Li & Lin (2024), extending them to accommodate the generalized algorithm and different smoothness assumptions.
Theorem 3.5 (Main, $\mathtt{Adam}$).
Let $\{L_{t}\}_{t=1}^{T}$ be independent stochastic losses satisfying 3.4 and that their expectation $L$ is ${\bm{H}}$-coordinate-wisely smooth w.r.t. $\ell_{\infty}$ norm. For $\mathtt{Adam}$ with $\beta_{1}=0$, we have that
$\min_{\frac{T}{2}<t\leq T}\mathbb{E}\left\|\nabla L({\bm{x}}_{t})\right\|_{1}$ $\leq O\left(E+\sqrt{E}\sqrt{\frac{\beta_{2}^{\frac{T}{4}}}{T(1-\beta_{2})}dv_{0}+\sum_{i=1}^{d}\sigma_{i}+d\sqrt{\epsilon}}\right)$
with
$E$ $=\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\left(1+\frac{\beta_{2}F}{T(1-\beta_{2})}\right)\left(\eta\sum_{i=1}^{d}H_{i}+\sqrt{1-\beta_{2}}\sum_{i=1}^{d}\sigma_{i}\right)$
and
$F=2\ln\left(1+\frac{\sum_{i=1}^{d}\sigma_{i}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\infty}^{2}+\max_{i\in[d]}H_{i}^{2}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}.$
We can determine the convergence rate of RMSprop by applying appropriate hyperparameters on Theorem 3.5. The optimal hyperparameters $\eta$ and $\beta_{2}$ can be selected by minimizing $E$. We would assume that $v_{0}+\epsilon>(\sum_{i=1}^{d}\sigma_{i}^{2}+\|\nabla L({\bm{x}}_{0})\|_{\infty}^{2}+\max_{i}H_{i}^{2}\eta^{2})/\mathtt{poly}(T)$ and $\frac{1}{1-\beta_{2}}=\mathtt{poly}(T)$. Then we can simplify the term by considering $F=O(\log T)$.
The two terms involving $\sum_{i=1}^{d}\sigma_{i}$ have a lower bound $\Theta\left(\sum_{i=1}^{d}\sigma_{i}\left(\frac{\log{T}}{T}\right)^{\frac{1}{2}}\right)$, which can reached by $1-\beta_{2}=\Theta\left(\frac{\log{T}}{T}\right)$. With this choice of $1-\beta_{2}$, the three terms involving $\eta$ has a lower bound $\Theta\left(\sqrt{\frac{\left(L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})\right)\sum_{i=1}^{d}H_{i}}{T}}\right)$ reached by $\eta=\Theta\left(\sqrt{\frac{L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})}{T\sum_{i=1}^{d}H_{i}}}\right)$. Such choices of hyperparameters can give the optimal convergence rate for stochastic case in Corollary 3.6. For convenience, we define $R\triangleq\left(L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})\right)\sum_{i=1}^{d}H_{i}$, which will be the core term in Corollaries 3.6 and 3.7.
Corollary 3.6 (Stochastic Case, general $\sigma_{i}$).
Let $\{L_{t}\}_{t=1}^{T}$ be independent stochastic losses satisfying 3.4 and that their expectation $L$ is ${\bm{H}}$-coordinate-wisely smooth w.r.t. $\ell_{\infty}$ norm. For $\beta_{1}=0$, $1-\beta_{2}=\Theta(\frac{\log{T}}{T})$, $\epsilon=0$, $\eta=\Theta\left(\sqrt{\frac{L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})}{T\sum_{i=1}^{d}H_{i}}}\right)$ and $v_{0}>(\sum_{i=1}^{d}\sigma_{i}^{2}+\|\nabla L({\bm{x}}_{0})\|_{\infty}^{2}+\max_{i}H_{i}^{2}\eta^{2})/\mathtt{poly}(T)$, we have that
$\min_{\frac{T}{2}<t\leq T}\mathbb{E}\left\|{\bm{g}}_{t}\right\|_{1}=O\left(\sqrt{\frac{R}{T}}+\sqrt{\sum_{i=1}^{d}\sigma_{i}}\left(\frac{R}{T}\right)^{\frac{1}{4}}+\sum_{i=1}^{d}\sigma_{i}\left(\frac{\log{T}}{T}\right)^{\frac{1}{4}}+\delta_{T}\right).$
with $\delta_{T}=\sqrt{\frac{dv_{0}}{T(1-\beta_{2})}}\exp{\left(-\frac{T(1-\beta_{2})}{8}\right)}\left[\left(\frac{R}{T}\right)^{\frac{1}{4}}+\sqrt{\sum_{i=1}^{d}\sigma_{i}}\left(\frac{\log{T}}{T}\right)^{\frac{1}{4}}\right]$
Here $\delta_{T}$ can be smaller than any polynomial of $T$ by manipulating the value of $\frac{T(1-\beta_{2})}{\log{T}}=\Theta(1)$. Then $\sum_{i=1}^{d}\sigma_{i}\left(\frac{\log{T}}{T}\right)^{\frac{1}{4}}$ is the leading term w.r.t. $T$ in the rate. The coefficient only involves $\sum_{i=1}^{d}\sigma_{i}$, suggesting that the rate can be much improved when noise is small. Below we get the convergence rate with the same hyperparameters in deterministic case in Corollary 3.7.
Corollary 3.7 (Deterministic Case, $\sigma_{i}=0$).
Let $\{L_{t}\}_{t=1}^{T}$ be deterministic losses satisfying 3.4 and that their expectation $L$ is ${\bm{H}}$-coordinate-wisely smooth w.r.t. $\ell_{\infty}$ norm. For $\beta_{1}=0$, $1-\beta_{2}=\Omega(\frac{\log{T}}{T})$, $\epsilon=0$, $\eta=\Theta\left(\sqrt{\frac{L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})}{T\sum_{i=1}^{d}H_{i}}}\right)$ and $v_{0}>(\sum_{i=1}^{d}\sigma_{i}^{2}+\|\nabla L({\bm{x}}_{0})\|_{\infty}^{2}+\max_{i}H_{i}^{2}\eta^{2})/\mathtt{poly}(T)$ for any polynomial $\mathtt{poly}(T)$, we have that
$\min_{\frac{T}{2}<t\leq T}\left\|{\bm{g}}_{t}\right\|_{1}=O\left(\sqrt{\frac{R}{T}}+\delta_{T}\right).$
with $\delta_{T}=\sqrt{\frac{dv_{0}}{T(1-\beta_{2})}}\exp{\left(-\frac{T(1-\beta_{2})}{8}\right)}\left(\frac{R}{T}\right)^{\frac{1}{4}}$.
Corollary 3.7 almost recovers Theorem 3.2, except the smoothness constant here $\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{(1,1)}$ is worse than that in Theorem 3.2, which is $\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{\infty}$, because it always holds that $\|\cdot\|_{1,1}\geq\|\cdot\|_{\infty}$. This gap is due to a technical difficulty of analyzing $\mathtt{Adam}$ or $\mathtt{RMSProp}$, as mentioned in the beginning of Section 3.2.
Dependence on $\epsilon$, $v_{0}$ and $\beta_{2}$.
While many previous works rely on the relatively large magnitude of $\epsilon$ compared to ${\bm{v}}_{t}$ and give a bound in the regime of $\mathtt{SGD}$ when the adaptive effect is dominated by the constant $\epsilon$ (Zaheer et al., 2018; De et al., 2018), our result actually prefers $\epsilon$ to be 00 while maintaining the value of $v_{0}+\epsilon$. We also note the dependence of our bound in Theorem 3.5 on $v_{0}$ is very mild and logarithmic. Theorem 3.5 has similar convergence rate for all $v_{0}$ of magnitude at most $\mathtt{poly}(T)$, while most previous result only addresses the case where $v_{0,i}$ is at the scale of noise (Li & Lin, 2024) or 00. The main reason for this adaptivity to a wide range of $v_{0}$ is our specific choice of $\beta_{2}=1-\Theta(\frac{\log T}{T})$, which allows the initial large $v_{0}$ to decay fast and resume normal training. Other existing results using $\beta_{2}=1-\Theta(1/T)$ (Défossez et al., 2022; Li & Lin, 2024) cannot allow large initial value $v_{0}$ because $v_{0}$ only decays a constant fraction throughout the training and the effective learning rate will be too small.
3.3 A unified analysis for blockwise $\mathtt{Adam}$
In this subsection, we present convergence analysis for a broader class of adaptive algorithms defined in Algorithm 3, which could be thought as a coarser version of $\mathtt{Adam}$. It does pre-conditioning blockwisely (specified by a partition function $\Phi:[d]\rightarrow[B]$ where $B$ is the number of blocks) instead of coordinate-wisely. Since $\mathtt{Adam}$ and $\mathtt{AdaSGD}$ can be viewed as special cases blockwise $\mathtt{Adam}$ (Algorithm 3) with $\Phi_{\mathtt{Adam}}:i\mapsto i$ and $\Phi_{\mathtt{AdaSGD}}:i\mapsto 1$ respectively, any convergence results for Algorithm 3 would imply convergence of $\mathtt{Adam}$ and $\mathtt{AdaSGD}$. Finally we also note that such blockwise $\mathtt{Adam}$ has been recently studied empirically by some concurrent work, where the algorithm is named by $\mathtt{Adamini}$ (Zhang et al., 2024b) and $\mathtt{Adalayer}$ (Zhao et al., 2024).
Algorithm 3 Blockwise $\mathtt{Adam}$
$\beta_{1},\beta_{2},\epsilon\geq 0$, block partition $\Phi:[d]\to[B]$, total steps $T$, learning rate schedule $\{\eta_{t}\}_{t=1}^{T}$, $\epsilon$, initial ${\bm{m}}_{0}$, $v_{0}$.
initialization ${\bm{x}}_{0}$, stochastic loss functions $\{L_{t}\}_{t=1}^{T}$
$v_{0,b}\leftarrow v_{0}$
for $t=1,2,\cdots,T$ :
   $g_{t,i}\leftarrow\nabla_{i}L_{t}({\bm{x}}_{t-1})$
   $m_{t,i}\leftarrow\beta_{1}m_{t-1,i}+(1-\beta_{1})g_{t,i}$
   $v_{t,b}\leftarrow\beta_{2}v_{t-1,b}+(1-\beta_{2})\left(\sum_{B(i)=b}g_{t,i}^{2}\right)/d_{b}$
   $x_{t,i}\leftarrow x_{t-1,i}-\eta_{t}\frac{m_{t,i}}{\sqrt{v_{t,B(i)}+\epsilon}}$
return ${\bm{x}}_{T}$
We first introduce some more notations. We define the vector ${\bm{x}}_{(b)}$ as $[x_{i}]_{\Phi(i)=b}$ and the submatrix $\mathbf{A}_{(b),(b^{\prime})}$ as $[A_{i,j}]_{\Phi(i)=b,\Phi(j)=b^{\prime}}$.
Definition 3.8 ($\Phi$-norm).
We define the $(\infty,2)$-norm w.r.t. partition $\Phi$ of vector ${\bm{x}}$ as the $\ell_{\infty}$ norm of the vector $\left(\frac{\left\|{\bm{x}}_{(b)}\right\|_{2}}{\sqrt{d_{b}}}\right)_{b=1}^{B}$, which is $\max_{b\in[B]}\frac{\left\|{\bm{x}}_{(b)}\right\|_{2}}{\sqrt{d_{b}}}$. For convenience, we will denote it by $\left\|{\bm{x}}\right\|_{\Phi}$ or just call it $\Phi$-norm. We denote its dual norm by $\|{\bm{x}}\|_{\Phi,*}$, which is equal to $\sum_{b=1}^{B}\sqrt{d_{b}}\left\|{\bm{x}}_{(b)}\right\|_{2}$.
Definition 3.9 (Generalized version of Definition 3.3).
For any partition function $\Phi:[d]\rightarrow[B]$ and $\mathbf{H}=(H_{1},\ldots,H_{B})\in\mathbb{R}^{B}$, we say a function $L$ is $\mathbf{H}$-smooth blockwisely w.r.t. $(\infty,2)$ norm , iff for any $b\in[B]$, ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d}$,
$\sqrt{d_{b}}\left\|\nabla_{(b)}L({\bm{x}})-\nabla_{(b)}L({\bm{y}})\right\|_{2}\leq H_{b}\left\|{\bm{x}}-{\bm{y}}\right\|_{\Phi}$
We further define the $\Phi$-smoothness of loss $L$ by $H(L,\Phi)=\sum_{b=1}^{B}H_{b}$, where $\{H_{b}\}_{b=1}^{B}$ are the smallest numbers making $L$ is $\mathbf{H}$-smooth blockwisely in the above sense.
We note that the above defined blockwise $(\infty,2)$-smoothness is both a generalization of the coordinate-wise smoothness defined in Definition 3.1 (corresponding to the case of each block only containing $1$ coordinate) and the standard $\ell_{2}$ smoothness (corresponding to the case of only having one block). In the former case, we have $B=d$, $\Phi_{\mathtt{Adam}}$ is the identity mapping $i\mapsto i$ and it holds that $H(L,i\mapsto i)\geq\sup_{{\bm{x}}\in\mathbb{R}^{d}}\left\|\nabla^{2}L({\bm{x}})\right\|_{1,1}\geq\sup_{{\bm{x}}\in\mathbb{R}^{d}}\left\|\nabla^{2}L({\bm{x}})\right\|_{\infty}$. In the latter case, we have $B=1$, $\Phi_{\mathtt{AdaSGD}}$ is the mapping $i\mapsto 1$ and $H(L,i\mapsto 1)=\sup_{{\bm{x}}\in\mathbb{R}^{d}}\left\|\nabla^{2}L({\bm{x}})\right\|_{2}$.
Similar to the coordinate-wise case, one can show that $\sum_{b=1}^{B}H_{b}$ w.r.t. partition $\Phi$ is an upper bound for the smoothness of loss $L$ w.r.t. to $\Phi$-norm.
Assumption 3.10 (Generalized version of 3.4).
There exists constant $\sigma_{b}$ such that
$\mathbb{E}\left\|\nabla_{(b)}L_{t}({\bm{x}})-\nabla_{(b)}L({\bm{x}})\right\|_{2}^{2}\leq d_{b}\sigma_{b}^{2}$
for any block $b\in[B],t\in\mathbb{N}$ and ${\bm{x}}\in\mathbb{R}^{d}$.
Theorem 3.11 (Main, Blockwise $\mathtt{Adam}$).
Under 3.10, suppose $L$ is $\mathbf{H}$-smooth blockwisely w.r.t. $(\infty,2)$-norm, where $\mathbf{H}=(H_{1},\ldots,H_{B})\in\mathbb{R}^{B}$, for Algorithm 3, we have that
$\min_{\frac{T}{2}<t\leq T}\mathbb{E}\sum_{b=1}^{B}\sqrt{d_{b}}\left\|\nabla_{(b)}L({\bm{x}}_{t})\right\|_{2}$ $\leq E+\sqrt{E}\sqrt{\frac{\beta_{2}^{\frac{T}{4}}}{T(1-\beta_{2})}d\sqrt{v_{0}}+\sum_{b=1}^{B}d_{b}\sigma_{b}+d\sqrt{\epsilon}}$
with
$E$ $=\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\left(1+\frac{\beta_{2}F}{T(1-\beta_{2})}\right)\left(\eta{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\sum_{b=1}^{B}H_{b}}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}\right),$
and
$F=2\ln\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}+\max_{b\in[B]}H_{b}^{2}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}.$
(1,1)-norm as a surrogate complexity measure for $H(L,\Phi_{\mathtt{Adam}})$.
$H_{i}$ in Definition 3.3 is determined by $\sup_{{\bm{x}}}\sum_{j=1}^{d}\left|\nabla^{2}_{i,j}L({\bm{x}})\right|$, which is difficult to compute because it requires taking supreme over the entire domain. An computationally-tractable alternative is to approximate $\sum_{i=1}^{d}H_{i}$ locally by the $(1,1)$-norm of Hessian of loss along the training trajectory. We provide an efficient approximation algorithm with guarantees by using hessian-vector product against random Cauchy vectors in Section D.2.
Different norms for smoothness.
As an implication of Theorem 3.11, we immediately get analogs of Corollaries 3.7, 3.7 and 3.6 for $\mathtt{AdaSGD}$, with the corresponding new noise assumption and smoothness assumption. When the optimization is not noise-dominated, i.e., the main term in the deterministic case $\sqrt{\frac{R}{T}}=\sqrt{\frac{(L({\bm{x}}_{0})-\min_{{\bm{x}}}L({\bm{x}}))H(L,\Phi)}{T}}$ becomes the leading term, the choice of $\Phi$ now matters a lot. Here the biggest change from $\mathtt{AdaSGD}$ to $\mathtt{Adam}$ is the difference between $H(L,\Phi_{\mathtt{AdaSGD}})$ and $H(L,\Phi_{\mathtt{Adam}})$, which roughly correspond to $\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{2}$ and $\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{1,1}$, using the above mentioned approximation.
Previous analyses of $\mathtt{Adam}$’s convergence (Shi & Li, 2021; Défossez et al., 2022; Li & Lin, 2024) usually assume smoothness under the $\ell_{2}$ norm. When using this assumption, the resulting convergence rate for $\mathtt{Adam}$ ends up being identical to the rate for $\mathtt{AdaSGD}$, which fails to capture why $\mathtt{Adam}$ often performs better than $\mathtt{AdaSGD}$ in practice. By shifting to an $\ell_{\infty}$ norm smoothness assumption, we can observe the key difference: the coefficient for $\mathtt{Adam}$’s convergence rate changes from $\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{2}$ to $\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{1,1}$, where the latter is typically much smaller when $\mathtt{Adam}$ optimizes faster. This change leads to a divergence in the rates of the two algorithms, and comparing these new coefficients can provide insight into which algorithm may be more effective under different conditions.
Finally, we note that the $\Phi_{\mathtt{Adam}}$-smoothness $H(L,\Phi_{\mathtt{Adam}})$ is not rotation-invariant in the sense that $H(L,\Phi_{\mathtt{Adam}})\neq H(L\circ\mathcal{R},\Phi_{\mathtt{Adam}})$ for a typical rotation $\mathcal{R}$. In practice, the $(1,1)$-norm of Hessian matrix can vary a lot when a rotation is performed on the loss as shown in Section 4.1. In contrast, $\Phi_{\mathtt{AdaSGD}}$-smoothness $H(L,\Phi_{\mathtt{AdaSGD}})$ is invariant under loss rotations.
3.4 Proof sketch of Theorem 3.11
We will use $\bar{{\bm{g}}}_{t}=\mathbb{E}[{\bm{g}}_{t}|{\bm{x}}_{<t}]=\nabla L({\bm{x}}_{t-1})$ to denote the full batch gradient. We first introduce Lemma 3.12, which is the main and only usage of our new smoothness assumption Definition 3.9.
Lemma 3.12.
For any twice differentiable loss which is $\mathbf{H}$-smooth block-wisely w.r.t. $(\infty,2)$-norm (Definition 3.9), we have for any ${\bm{x}}$ and ${\bm{\Delta}}\in\mathbb{R}^{d}$,
${\bm{\Delta}}^{\top}\nabla^{2}L({\bm{x}}){\bm{\Delta}}\leq\sum_{b=1}^{B}\frac{H_{b}}{d_{b}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}.$ (1)
We start by considering the decrease of $L({\bm{x}}_{t})$ in a single step $t$. We can upper bound the second order term in the Taylor expansion (Equation 2) with Lemma 3.12. Then we can get loss decrease in a single step
$L({\bm{x}}_{t})-L({\bm{x}}_{t-1})$ $\leq-\eta\sum_{i=1}^{d}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,\Phi(i)}+\epsilon}}+\frac{1}{2}\eta^{2}\sum_{b=1}^{B}\frac{H_{b}}{d_{b}}\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}}{v_{t,b}+\epsilon}$ (2)
$\leq-\eta\sum_{b=1}^{B}\frac{{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}+\frac{1}{2}\eta^{2}\sum_{b=1}^{B}H_{b}\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}{v_{t,b}+\epsilon}$ (3)
The proof contains two main parts: lower bounding the first order term and upper bounding the second order term. Below we address the second-order term first. At each step, the second order term $\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}{v_{t,b}+\epsilon}$ can be as large as $\frac{1}{1-\beta_{2}}$. But if we sum over all the steps, we can employ Lemma 3.13 to bound the sum by $T+\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}$ rather than $\frac{T}{1-\beta_{2}}$.
Lemma 3.13.
Given any $10<\beta_{2}<$ 1, for any scalar sequences $\{v_{t}\}_{t=0}^{T}$ and $\{g_{t}\}_{t=1}^{T}$ satisfy that $v_{0}\geq 0,v_{1}>0$ and $v_{t}-\beta_{2}v_{t-1}\geq(1-\beta_{2})g_{t}^{2}$ for $t\geq 1$, it holds that
$\sum_{t=1}^{T}\frac{g_{t}^{2}}{v_{t}}\leq T+\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{v_{T}}{v_{0}}}.$ (4)
Now we turn to the first term. Ideally, we would like to connect the first order term to $\|\bar{\bm{g}}_{t,(b)}\|_{2}^{2}$ by taking expectation, e.g., $\mathbb{E}_{t}\frac{{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}\approx\frac{\mathbb{E}_{t}{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}=\frac{\|\bar{\bm{g}}_{t,(b)}\|_{2}^{2}}{\sqrt{v_{t,b}+\epsilon}}$, where we use $\mathbb{E}_{t}[\cdot]$ as abbreviation for $\mathbb{E}[\cdot|{\bm{x}}_{<t}]$. However, this is not correct because both the numerator and denominator in $\sum_{b=1}^{B}\frac{{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}$ depend on the stochastic gradient noise at step $t$. To circumvent this difficulty, we use Lemma 3.14 to lower bound each $\mathbb{E}_{t}\frac{{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}$ by $\frac{\mathbb{E}_{t}{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{\mathbb{E}_{t}v_{t,b}+\epsilon}}\geq\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}$ minus error terms related to noise magnitude $\sigma_{b}$, where $\tilde{v}_{t,b}=\beta_{2}v_{t-1,b}+(1-\beta_{2})\left(\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}\right)$.
Lemma 3.14 (first-order approximation).
For any block $b\in[B]$, it holds that
$\mathbb{E}\sum_{t=1}^{T}\frac{{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}$ $\geq\frac{1}{2}\mathbb{E}\sum_{t=1}^{T}\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}-\sqrt{1-\beta_{2}}Td_{b}\sigma_{b}-\frac{d_{b}\sigma_{b}\beta_{2}}{\sqrt{1-\beta_{2}}}\mathbb{E}\left[\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}\right].$ (5)
Finally, we employ Cauchy inequality to upper bound $\left\|\bar{{\bm{g}}}_{t}\right\|_{\Phi,*}$:
$\sum_{t=\frac{T}{2}+1}^{T}\left\|\bar{{\bm{g}}}_{t}\right\|_{\Phi,*}=\sum_{t=\frac{T}{2}+1}^{T}\sum_{b=1}^{B}\sqrt{d_{b}}\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}\leq\sqrt{\sum_{t=\frac{T}{2}+1}^{T}\sum_{b=1}^{B}\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}}\sqrt{\sum_{t=\frac{T}{2}+1}^{T}\sum_{b=1}^{B}d_{b}\sqrt{\tilde{v}_{t,b}+\epsilon}}.$ (6)
We use Lemma 3.15 to upper bound $\sum_{t=\frac{T}{2}+1}^{T}d_{b}\sqrt{\tilde{v}_{t,b}+\epsilon}$ by $\sum_{t=1}^{T}\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}$. $\sum_{t=1}^{T}\sum_{b=1}^{B}\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}$ can be bounded from analysis above, which finishes the proof.
Lemma 3.15.
For any block $b\in[B]$, it holds that
$\sum_{t=\frac{T}{2}+1}^{T}\mathbb{E}\left[\sqrt{\tilde{v}_{t,b}+\epsilon}\right]$ $\leq\frac{2\beta_{2}^{\frac{T}{4}}}{1-\beta_{2}}\sqrt{v_{0,b}}+\frac{T}{2}\sigma_{b}+\frac{T}{2}\sqrt{\epsilon}+2\sum_{t=1}^{T}\mathbb{E}\left[\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\right].$ (7)
4 Experiments
In order to empirically investigate and confirm the implications of our propsosed theory, we compare the training performance of $\mathtt{Adam}$ with $\mathtt{AdaSGD}$, $\mathtt{SGD}$ and $\mathtt{rotated\ Adam}$ on multiple different tasks. In $\mathtt{Adam}$ and its variants (including AdaSGD) we always set $\beta_{1},\beta_{2}=(0.9,0.99)$. Momentum in SGD is also set to $0.9$. Weight decay is always deactivated.
4.1 Quadratic loss
We perform controlled experiments on quadratic loss to study the relationship between optimization speed of $\mathtt{Adam}$ and the shape of Hessian, in terms of $\Phi_{\mathtt{Adam}}$-smoothness. More specifically, we consider $\Sigma=\text{diag}(\underbrace{1,\cdots,1}_{10},1,\frac{1}{2^{2}},\frac{1}{3^{2}},\cdots,\frac{1}{990^{2}})\in\mathbb{R}^{1000\times 1000}$ and optimize the corresponding quadratic loss of $\Sigma$ by $\mathtt{Adam}$, with different levels of rotations. We manually generate orthogonal matrices $\mathcal{R}_{i}$ in the following way. We first sample $\mathcal{M}\in\mathbb{R}^{d\times d}$ where $\mathcal{M}_{i,j}$ is i.i.d. sampled from $N(0,1)$. Then $\mathcal{A}=\mathcal{M}-\mathcal{M}^{\top}$ is a skew-symmetric matrix and $\exp{(t\mathcal{A})}$ represents a continuous family of matrices. We define $\mathcal{R}_{i}=\exp{(t_{i}\mathcal{A})}$ for different $t_{i}$. When $t_{i}=0$, we know $\mathcal{R}_{i}=I$. When $t_{i}\to\infty$, $\mathcal{R}_{i}$ converges to a random orthogonal matrix in distribution.
Then we optimize $L_{0}({\bm{x}})={\bm{x}}^{\top}\Sigma{\bm{x}}$ with $\mathtt{AdaSGD}$ and $\mathtt{Adam}$ and optimize $L_{i}({\bm{x}})={\bm{x}}^{\top}\mathcal{R}_{i}^{\top}\Sigma\mathcal{R}_{i}{\bm{x}}$ with $\mathtt{Adam}$ for $100$ steps. Because $\mathtt{AdaSGD}$ is rotational invariant, the optimization performance of $\mathtt{AdaSGD}$ is the same on all $L_{i}$. We tune learning rates for each setting and present their best results in Table 1.
We find a clear pattern that $\mathtt{Adam}$ optimizes worse when the $(1,1)$ norm of Hessian matrix increases, as suggested by our Corollary 3.7. Moreover, when $(1,1)$ norm divided by dimension is smaller than spectral norm, $\mathtt{Adam}$ tends to optimize faster than $\mathtt{AdaSGD}$ and vice versa, as suggested by our Theorem 3.11.
Optimizer $(1,1)$-norm$/d$ Loss ($\beta_{1}=\beta_{2}=0$) Loss ($\beta_{1}=0.9,\beta_{2}=0.99$)
AdaSGD $0.00582$ $0.00881$ $0.00172$
Adam $0.00582$ $0.00030$ $0.00001$
Adam ($\mathcal{R}_{1}$) $0.04162$ $0.00317$ $0.00062$
Adam ($\mathcal{R}_{2}$) $0.25364$ $0.00588$ $0.00122$
Adam ($\mathcal{R}_{3}$) $0.61866$ $0.00747$ $0.00179$
Adam ($\mathcal{R}_{4}$) $1.29959$ $0.00920$ $0.00239$
Table 1: The final loss values obtained by different optimizers and the $(1,1)$-norm of Hessian matrix for the corresponding unrotated objective and rotated objectives. The spectral norm of the Hessian matrix is always $1$. $\mathtt{Adam}$ optimizes worse when the $(1,1)$-norm of Hessian matrix increases, as suggested by our Corollary 3.7. Moreover, when $(1,1)$-norm is smaller than spectral norm times space dimension, $\mathtt{Adam}$ tends to optimize faster than $\mathtt{AdaSGD}$ and vice versa, which justifies the effectiveness of $\Phi$-smoothness as a tool to predict the optimization speed of blockwise $\mathtt{Adam}$.
4.2 GPT-2 on language modeling task
We train GPT-2 small (124M parameters) [^3] We use nanoGPT codebase, https://github.com/karpathy/nanoGPT on the OpenWebText corpus containing more than 9B tokens for 100k iterations with sequence length of $512$ sequence length and $480$ sentences per batch. We use cosine learning rate schedule of the same peak learning rate $6\times 10^{-4}$, which is also the default of nanoGPT codebase. We did a grid search to find the maximum possible peak learning rate for $\mathtt{SGD}$  [^4] We tried $0.00001,0.00003,0.0001,0.0003,0.001,0.003,0.01,0.03,0.1,0.3,1$. . The training losses and evaluation losses of different optimizers are plotted in Figure 1. As mentioned in Section 1, $\mathtt{Adam}$ converges faster than $\mathtt{AdaSGD}$ while they both converge faster than $\mathtt{rotated\ Adam}$. Since we propose the $(1,1)$-norm of Hessian as a non rotation-invariant metric that can affect the convergence rate of $\mathtt{Adam}$, we also measure it for original loss function $L$ and rotated loss function $\tilde{L}$ on checkpoints trained with different losses. The results are presented in Table 2. The same correlation between norms and convergence rates holds here. The smaller the norm is, the faster the optimizer works.
Optimizer Smoothness Metric Matrix Norm Estimated Value
AdaSGD $H(L,\Phi_{\mathtt{AdaSGD}})$ $\left\|\nabla^{2}L({\bm{x}})\right\|_{2}$ $24.86$
Adam $H(L,\Phi_{\mathtt{Adam}})$ $\left\|\nabla^{2}L({\bm{x}})\right\|_{1,1}$ $3.2$
Rotated Adam $H(L\circ\mathcal{R},\Phi_{\mathtt{Adam}})$ $\left\|R^{\top}\nabla^{2}L({\bm{x}})R\right\|_{1,1}$ $36.16$
Table 2: Hessian norms for the last GPT-2 checkpoints trained with different optimizers.
GPT-2 small models have more than $100$ million parameters, and thus the size of its hessian of loss as well as the rotation matrix is more than $10^{16}$, which is way more than the storage of the modern computers. Thus our experiments face the following two implementation challenges. Below we briefly discuss them and leave the full details in Appendix D.
Efficient generation of random orthogonal matrices.
Due to the huge parameter count, it is computationally infeasible to sample a random orthogonal matrix from Haar measure over the orthogonal group. Thus we decide to sample from an alternative distribution over orthogonal matrices, where the final rotation matrix is the composition of a sequence of simple orthogonal matrices, including random permutation and apply random orthogonal transformation on the both side of matrix-valued parameters. See details in Section D.1.
Efficient estimation of $(1,1)$-norm of Hessian matrix.
We first subsample a fixed batch of training data for estimating the $(1,1)$-norm of Hessian matrix. The high-level idea here is to compute the Hessian vector product between Hessian of training loss on this batch and a sequence of random Cauchy vectors and take the $\ell_{1}$ norm of the coordinate-wise median of the resulting sequence of Hessian vector products. Because Cauchy distribution is 1-stable, the resulting product is also a vector of Cauchy random variables, and each of them has the magnitude equal to $\ell_{1}$ norm of the corresponding row of the Hessian. Thus when we have infinitely many samples, the $\ell_{1}$ norm of the coordinate-wise median converges almost surely to the $(1,1)$-norm of Hessian. See details in Section D.2, where we also provide a non-asymptotic high-probability multiplicative bound for the estimation error.
4.3 ResNet18 on CIFAR-10
To further test the correlation between $\Phi$-smoothness and the optimization performance holds for architectures other than transformers, we conduct a similar experiment on ResNet18 trained on CIFAR-10 (Krizhevsky & Hinton, 2009). We applied random crop and random horizontal flip augmentations over the training data to promote better generalization. We tuned each optimizer through searching over the same grid of learning rates [^5] We used the following values: $6.25\times 10^{-4}$, $1.25\times 10^{-3}$, $2.5\times 10^{-3}$, $5.0\times 10^{-3}$, $1.0\times 10^{-2}$, $2.0\times 10^{-2}$, $4.0\times 10^{-2}$, $8.0\times 10^{-2}$, $1.6\times 10^{-1}$, $3.2\times 10^{-1}$, $6.4\times 10^{-1}$, $1.28\times 10^{0}$. The number of iterations is adjusted per batch size to result in  20 epochs for each training run (for instance, 4000 iterations were used for a batch size of 256, and 1000 iterations were used for a batch size of 1024). For the training loss and training accuracy plotted in Figure 2, it is measured on a subset of augmented training data that is the same size of evaluation set. The evaluation loss and accuracy are measured on the entire evaluation set without the augmentation. Track running stats is set to false at initialization.
Figure 2: Training and test losses (left) and accuracy (right) of ResNet18 on CIFAR-10 with $\mathtt{Adam}$, $\mathtt{AdaSGD}$, $\mathtt{rotated\ Adam}$, and $\mathtt{SGD}$. We use batch size $256$ and the optimal learning rate in terms of training loss from grid search. Solid and dashed lines correspond to the training and evaluation set metrics respectively. $\mathtt{Adam}$ converges faster than other algorithms.
Figure 2 depicts the loss and accuracy curves for the best performing hyperparameters chosen over the training set’s final loss for batch size $256$. [^6] We have intentionally limited the number of training iterations to emphasize the difference of optimizers in terms of training speed over generalization. We also provide the results for other choices of batch size in Table 3. When it comes to optimization speed, even for ResNet18, $\mathtt{Adam}$ is always better than and they are always better than $\mathtt{AdaSGD}$, $\mathtt{SGD}$, and $\mathtt{rotated\ Adam}$ across different batch sizes. Note that this does not contradict with common practice of training ResNet with $\mathtt{SGD}$, where the main goal is to get better generalization and the training budget is large so all optimizers can easily achieve full training accuracy. In our experiment, we study optimization speed and intentionally limit the number of steps.
Batch Size SGD AdaSGD Adam Rotated Adam
$16$ $0.0777$ $0.114$ $0.064$ $0.0905$
$64$ $0.0698$ $0.0854$ $0.0472$ $0.0574$
$256$ $0.0723$ $0.0787$ $0.0359$ $0.0485$
$1024$ $0.1115$ $0.0915$ $0.0735$ $0.0817$
Table 3: Training losses of ResNet for different optimizers and different batch sizes within $20$ epochs. For each setting, we choose the optimal performance over all the learning rates. The performance of $\mathtt{Adam}$ is consistently the best among all four optimizers.
We also measure the Hessian for checkpoints obtained at batch size $256$ and the results are in Table 4. The correlation between norms and convergence rates still holds here. When the $(1,1)$-norm divided by $d$ is smaller than spectral norm, $\mathtt{Adam}$ optimizes faster than $\mathtt{AdaSGD}$.
Optimizer Smoothness Metric Matrix Norm Estimated Value
AdaSGD $H(L,\Phi_{\mathtt{AdaSGD}})$ $\left\|\nabla^{2}L({\bm{x}})\right\|_{2}$ $1.5355$
Adam $H(L,\Phi_{\mathtt{Adam}})$ $\left\|\nabla^{2}L({\bm{x}})\right\|_{1,1}$ $0.0036$
Rotated Adam $H(L\circ\mathcal{R},\Phi_{\mathtt{Adam}})$ $\left\|R^{\top}\nabla^{2}L({\bm{x}})R\right\|_{1,1}$ $0.9868$
Table 4: Hessian norms for optimal ResNet checkpoints trained with different optimizers and batch size $256$.
5 Related Works
Comparison between Adam and SGD
Previous work tries to analyze the difference between $\mathtt{Adam}$ and $\mathtt{SGD}$ from different perspectives. Zhou et al. (2018) proves a faster convergence rate of $\mathtt{Adam}$ than $\mathtt{SGD}$ when the stochastic gradients are sparse. Zhang et al. (2020) suggests that $\mathtt{SGD}$ suffers more from heavy-tailed noise than $\mathtt{Adam}$. Pan & Li (2023) claims that $\mathtt{Adam}$ has lower directional sharpness because of the effect of coordinate-wise clipping. Other works also consider the coordinate-wise normalization of $\mathtt{Adam}$ (Balles & Hennig, 2018; Kunstner et al., 2022). Kunstner et al. (2024) shows that the heavy-tailed class imbalance in language modeling tasks will cause $\mathtt{SGD}$ to converge slower when it can only optimize majority class well. Zhang et al. (2024a) finds that $\mathtt{Adam}$ is better at handling the block heterogeneity of Hessian matrix, which is a specific phenomenon in transformers. When viewing $\mathtt{Adam}$ as an adaptive method, there are works showing that adaptive methods have an advantage of achieving optimal convergence rate without relying on problem-dependent constant (Ward et al., 2020; Levy et al., 2021).
Convergence rate of Adam
There are many works showing convergence rate for $\mathtt{Adam}$ (Zhou et al., 2018; Chen et al., 2018; Zou et al., 2019; Shi & Li, 2021; Guo et al., 2021; Défossez et al., 2022; Zhang et al., 2022b). Most of them rely on the smoothness of the loss function, which is measured w.r.t. $\ell_{2}$ norm. Zhang et al. (2019) proposes the $(L_{0},L_{1})$ smoothness condition should be more reasonable than globally bounded smoothness. Li et al. (2024) further generalizes the $(L_{0},L_{1})$ smoothness condition. However, they still focus on the default $\ell_{2}$ norm which is rotation-invariant. To the best of our knowledge, we are the first to assume gradient Lipschitzness under $\ell_{\infty}$ norm for the analysis on $\mathtt{Adam}$.
Comparison with Li & Lin (2024)
Li & Lin (2024) employs the same $\ell_{1}$ norm for gradient and improves the dependence on dimension $d$ compared to previous results for $\ell_{2}$ norm. But they still assume the common $\ell_{2}$ norm smoothness while we adapt their results under $\ell_{\infty}$ norm smoothness to potentially further improve dependence on $d$. Another drawback of Li & Lin (2024) is setting ${\bm{v}}_{0}$ based on noise magnitude $\sigma$. which is impractical in real experiments because $\sigma$ is unknown. Overestimation for $\sigma$ will result in slow convergence because large ${\bm{v}}_{0}$ causes $\mathtt{Adam}$ to behave similarly with $\mathtt{SGD}$ without adjusting the coordinate-wise learning rate adaptively. In contrast, we allow for general initialization for ${\bm{v}}_{0}$ and our convergence rate can work well in both noisy setting and deterministic setting. We also use $1-\beta_{2}=\Theta\left(\frac{\log{T}}{T}\right)$ to obtain our convergence rate while Li & Lin (2024) requires $1-\beta_{2}=\Theta\left(\frac{1}{T}\right)$.
6 Conclusion
We give a new convergence analysis (Theorem 3.5) for $\mathtt{Adam}$ in the stochastic non-convex setting using a novel smoothness assumption. We show the convergence rate for the 1-norm of the gradient is $O\left(\frac{1}{\sqrt{T}}\right)$ in the deterministic case (Corollaries 3.7 and 3.7) and $O\left(\left(\frac{\log T}{T}\right)^{1/4}\right)$ in the stochastic case (Corollary 3.6). We also extend our analysis to blockwise $\mathtt{Adam}$ on loss $L$ with respect to an arbitrary partition of the parameters $\Phi$ (Theorem 3.11) using the corresponding smoothness $H(L,\Phi)$ (Definition 3.9). Our bound for $\mathtt{Adam}$ involves $(1,1)$-norm of Hessian, rather than the spectral norm of Hessian, which is relevant to the convergence speed of $\mathtt{AdaSGD}$. This leads to significantly better smoothness conditions for deep learning models including ResNet-18 and GPT2 empirically. Our experiments also verify that the smoothness measure $H(L,\Phi)$ positively correlates with the optimization speed of blockwise $\mathtt{Adam}$ with respect to the partition $\Phi$.
Acknowledgement
The authors would like to thank Khashayar Gatmiry and Sharut Gupta for helpful discussion and preliminary experiments in exploring the idea of rotated Adam. ZL is supported by OpenAI.
Appendix A Convergence rate of $\mathtt{SignGD}$ for deterministic loss
Proof of Theorem 3.2.
We will directly prove a more general verion of Theorem 3.2. Because $L$ is $H$-smooth with respect to $\|\cdot\|_{\infty}$, we have that
$L({\bm{x}}_{t+1})-L({\bm{x}}_{t})$ $\leq-\nabla L({\bm{x}}_{t})^{\top}({\bm{x}}_{t}-{\bm{x}}_{t+1})+\frac{H}{2}\left\|{\bm{x}}_{t}-{\bm{x}}_{t+1}\right\|^{2}$
$\leq-\eta\left\|\nabla L({\bm{x}}_{t})\right\|_{*}+\frac{\eta^{2}H}{2}\eta^{2}$ (8)
This implies that
$\min_{1\leq t\leq T}\left\|\nabla L({\bm{x}}_{t})\right\|_{*}\leq\frac{1}{T}\sum_{t=1}^{T}\left\|\nabla L({\bm{x}}_{t})\right\|_{*}\leq\frac{L({\bm{x}}_{0})-L({\bm{x}}_{T})}{T\eta}+\frac{H\eta}{2},$
which completes the proof. ∎
Appendix B Invariance property of $\mathtt{Adam}$ and $\mathtt{SGD}$
See 2.2
Proof of Theorem 2.2.
For $\mathtt{SGD}$ and $\mathtt{AdaSGD}$, we will show they are rotation-invariant by induction. For any rotating transformation $\mathcal{R}({\bm{x}})=\bm{R}{\bm{x}}$, suppose $\tilde{{\bm{x}}}_{s}=\mathcal{R}^{-1}({\bm{x}}_{s})=\bm{R}^{\top}{\bm{x}}_{s}$ holds for $s\leq t-1$. Then we have that $\tilde{{\bm{g}}}_{t}=\nabla_{\tilde{{\bm{x}}}}\tilde{L}_{t}(\tilde{{\bm{x}}}_{t})=\bm{R}^{\top}\nabla_{{\bm{x}}}L(\bm{R}^{-1}\tilde{{\bm{x}}}_{t-1})=\bm{R}^{\top}\nabla_{{\bm{x}}}L({\bm{x}}_{t-1})=\bm{R}^{\top}{\bm{g}}_{t}$ and $\tilde{{\bm{m}}}_{t}=\bm{R}^{\top}{\bm{m}}_{t}$. From the update rule of $\mathtt{SGD}$, we have that $\tilde{{\bm{x}}}_{t}=\tilde{{\bm{x}}}_{t-1}-\eta_{t}\tilde{{\bm{m}}}_{t}=\bm{R}^{\top}{\bm{x}}_{t-1}-\eta_{t}\bm{R}^{\top}{\bm{m}}_{t}=\bm{R}^{\top}({\bm{x}}_{t-1}-\eta_{t}{\bm{m}}_{t})=\bm{R}^{\top}{\bm{x}}_{t}$. For the update rule of $\mathtt{AdaSGD}$, we further have that $\left\|\tilde{{\bm{g}}}_{t}\right\|_{2}^{2}=\left\|{\bm{g}}_{t}\right\|_{2}^{2}$ because $\bm{R}$ is an orthogonal matrix. Then $\tilde{v}_{t}=v_{t}$ and the derivation is similar.
For $\mathtt{Adam}$ and $\mathtt{SignGD}$, it is easy to show by induction they are invariant w.r.t. any permutating transformation because the operation on gradient is performed on each coordinate separately. We only need to show they are not invariant w.r.t. a rotating transformation. We choose $\bm{R}=[\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}};\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}}]$, $L_{t}({\bm{x}})=L({\bm{x}})=2x_{1}^{2}+x_{2}^{2}$. Due to the update rule of $\mathtt{SignGD}$, it can only update ${\bm{x}}$ and $\tilde{{\bm{x}}}$ in the direction of $[1,1]$ and $[1,-1]$. But when rotating the update direction on $\tilde{{\bm{x}}}$ back to the space of ${\bm{x}}$. The update direction can only be $[1,0]$ or $[0,1]$ that are different from the update direction in the original space. Because the first step in $\mathtt{Adam}$ takes the same direction in $\mathtt{SignGD}$, we simultaneously show that both $\mathtt{SignGD}$ and $\mathtt{Adam}$ are not rotation-invariant. ∎
Appendix C Proof Details
C.1 Proof for convergence rate of blockwise $\mathtt{Adam}$
We also need new notations and different assumptions for specific partition $\Phi(\cdot)$. For specific block $b$, $(b)$ is defined as $\Phi^{-1}(b)=\{i|\Phi(i)=b\}$ and $d_{b}=\#(b)$, the number of parameters in block $b$.
As mentioned in Section 3.4, we will use Lemma 3.12 to bound the second order term under smoothness Definition 3.9 and use Lemma 3.13 to better control the growth of the sum of second order term. See 3.12
Proof of Lemma 3.12.
From Definition 3.9, we know that
$H_{b}$ $\geq\sup_{{\bm{x}},{\bm{\Delta}}}\frac{\sqrt{d_{b}}\left\|\nabla_{(b)}L({\bm{x}}+{\bm{\Delta}})-\nabla_{(b)}L({\bm{x}})\right\|_{2}}{\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}}}$
$=\sup_{{\bm{x}},{\bm{\Delta}}}\frac{\sqrt{d_{b}}\left\|\nabla_{(b),:}^{2}L({\bm{x}}){\bm{\Delta}}\right\|_{2}}{\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}}}$
$=\sup_{{\bm{x}},{\bm{\Delta}}}\frac{\sqrt{d_{b}}\left\|\sum_{b^{\prime}=1}^{B}\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}}}$
$=\sup_{{\bm{x}},{\bm{\Delta}},\left\|{\bm{\Delta}}^{\prime}_{(b)}\right\|_{2}\leq 1}\frac{\sqrt{d_{b}}\left\langle{\bm{\Delta}}^{\prime}_{(b)},\sum_{b^{\prime}=1}^{B}\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle}{\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}}}$
$=\sup_{{\bm{x}},\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}\leq\sqrt{d_{b^{\prime}}},\left\|{\bm{\Delta}}^{\prime}_{(b)}\right\|_{2}\leq 1}\sqrt{d_{b}}\left\langle{\bm{\Delta}}^{\prime}_{(b)},\sum_{b^{\prime}=1}^{B}\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle$
$=\sup_{{\bm{x}},{\bm{\Delta}},{\bm{\Delta}}^{\prime}}\sum_{b^{\prime}=1}^{B}\frac{\sqrt{d_{b}}\sqrt{d_{b^{\prime}}}}{\left\|{\bm{\Delta}}^{\prime}_{(b)}\right\|_{2}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}\left\langle{\bm{\Delta}}^{\prime}_{(b)},\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle$
Then for any ${\bm{x}}$ and ${\bm{\Delta}}$, we know that
$\frac{H_{b}}{d_{b}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}$ $\geq\frac{\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}}{d_{b}}\sum_{b^{\prime}=1}^{B}\frac{\sqrt{d_{b}}\sqrt{d_{b^{\prime}}}}{\left\|{\bm{\Delta}}_{(b)}\right\|_{2}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}\left\langle{\bm{\Delta}}_{(b)},\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle$
$=\sum_{b^{\prime}=1}^{B}\frac{\sqrt{d_{b^{\prime}}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}}{\sqrt{d_{b}}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}\left\langle{\bm{\Delta}}_{(b)},\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle$
and
$2\sum_{b=1}^{B}\frac{H_{b}}{d_{b}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}$
$=$ $\sum_{b=1}^{B}\frac{H_{b}}{d_{b}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}+\sum_{b^{\prime}=1}^{B}\frac{H_{b^{\prime}}}{d_{b^{\prime}}}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}^{2}$
$\geq$ $\sum_{b=1}^{B}\sum_{b^{\prime}=1}^{B}\frac{\sqrt{d_{b^{\prime}}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}}{\sqrt{d_{b}}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}\left\langle{\bm{\Delta}}_{(b)},\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle+\sum_{b^{\prime}=1}^{B}\sum_{b=1}^{B}\frac{\sqrt{d_{b}}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}}\left\langle{\bm{\Delta}}_{(b^{\prime})},\nabla_{(b^{\prime}),(b)}^{2}L({\bm{x}}){\bm{\Delta}}_{(b)}\right\rangle$
$\geq$ $2\sum_{b=1}^{B}\sum_{b^{\prime}=1}^{B}{\bm{\Delta}}_{(b)}^{\top}\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}=2{\bm{\Delta}}^{\top}\nabla^{2}L({\bm{x}}){\bm{\Delta}}.$
The last inequality comes from mean inequality. ∎
See 3.13
Proof of Lemma 3.13.
Notice that $1-x\leq\ln{\frac{1}{x}}$ for any positive $x$. We can have that
$\sum_{t=1}^{T}\frac{g_{t}^{2}}{v_{t}}$ $\leq\sum_{t=1}^{T}\frac{v_{t}-\beta_{2}v_{t-1}}{(1-\beta_{2})v_{t}}$
$=\sum_{t=1}^{T}\left[1+\frac{\beta_{2}}{1-\beta_{2}}\left(1-\frac{v_{t-1}}{v_{t}}\right)\right]$
$\leq T+\frac{\beta_{2}}{1-\beta_{2}}\sum_{t=1}^{T}\ln{\frac{v_{t}}{v_{t-1}}}$
$=T+\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{v_{T}}{v_{0}}}.$ (9)
when $v_{0}\neq 0$. When $v_{0}=0$, we can still have that
$\sum_{t=1}^{T}\frac{g_{t}^{2}}{v_{t}}$ $\leq\frac{1}{1-\beta_{2}}+\sum_{t=2}^{T}\frac{g_{t}^{2}}{v_{t}}$
$\leq\frac{1}{1-\beta_{2}}+(T-1)+\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{v_{T}}{v_{1}}}$
$=T+\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{v_{T}}{v_{1}/e}}.$
Next we deal with the first order term by approximating it with a deterministic term. Here we need some new notations. Recall that ${\bm{g}}_{t}$ denotes the gradient of mini-batch $L_{t}({\bm{x}}_{t-1})$ at step $t$. And $\mathbb{E}\left[{\bm{g}}_{t}\middle|{\bm{x}}_{t-1}\right]=\nabla L({\bm{x}}_{t-1})$ because $\mathbb{E}{L_{t}}=L$. The full-batch gradient is $\bar{{\bm{g}}}_{t}=\nabla L({\bm{x}}_{t-1})$. Different kinds of second-order momentum are defined in the following way.
$v_{t,b}$ $=\beta_{2}^{t}\left\|{\bm{g}}_{1,(b)}\right\|_{2}^{2}/d_{b}+\left(1-\beta_{2}\right)\sum_{j=0}^{t-1}\beta_{2}^{j}\left(\left\|{\bm{g}}_{t-j,(b)}\right\|_{2}^{2}\right)/d_{b}$
$\tilde{v}_{t,b}$ $=(1-\beta_{2})\left(\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}\right)+\beta_{2}v_{t-1,b}$
See 3.14
Proof of Lemma 3.14.
The first order change in block $b$ can decomposed into two terms.
$\mathbb{E}\sum_{t=1}^{T}\sum_{\Phi(i)=b}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,b}+\epsilon}}$ $=\mathbb{E}\sum_{t=1}^{T}\sum_{\Phi(i)=b}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}+\mathbb{E}\left[\sum_{t=1}^{T}\sum_{\Phi(i)=b}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,b}+\epsilon}}-\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\right]$ (10)
$=\mathbb{E}\sum_{t=1}^{T}\sum_{\Phi(i)=b}\mathbb{E}\left[\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\middle|{\bm{x}}_{t-1}\right]+\mathbb{E}\left[\sum_{t=1}^{T}\sum_{\Phi(i)=b}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,b}+\epsilon}}-\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\right]$
$=\mathbb{E}\sum_{t=1}^{T}\sum_{\Phi(i)=b}\frac{\bar{g}_{t,i}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}+\mathbb{E}\left[\sum_{t=1}^{T}\sum_{\Phi(i)=b}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,b}+\epsilon}}-\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\right]$
For the second term, we have that
$\sum_{\Phi(i)=b}\left|g_{t,i}\bar{g}_{t,i}\left(\frac{1}{\sqrt{v_{t,b}+\epsilon}}-\frac{1}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\right)\right|$
$=$ $\sum_{\Phi(i)=b}\frac{\left|g_{t,i}\bar{g}_{t,i}\left(\tilde{v}_{t,b}-v_{t,b}\right)\right|}{\sqrt{v_{t,b}+\epsilon}\sqrt{\tilde{v}_{t,b}+\epsilon}\left(\sqrt{v_{t,b}+\epsilon}+\sqrt{\tilde{v}_{t,b}+\epsilon}\right)}$
$=$ $\sum_{\Phi(i)=b}\frac{\left|g_{t,i}\bar{g}_{t,i}(1-\beta_{2})\left(\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}-\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}\right)\right|}{\sqrt{v_{t,b}+\epsilon}\sqrt{\tilde{v}_{t,b}+\epsilon}\left(\sqrt{v_{t,b}+\epsilon}+\sqrt{\tilde{v}_{t,b}+\epsilon}\right)}$
$=$ $\sum_{\Phi(i)=b}\frac{\left|g_{t,i}\bar{g}_{t,i}(1-\beta_{2})\left(\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}}+\sqrt{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}\right)\left(\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}}-\sqrt{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}\right)\right|}{\sqrt{v_{t,b}+\epsilon}\sqrt{\tilde{v}_{t,b}+\epsilon}\left(\sqrt{v_{t,b}+\epsilon}+\sqrt{\tilde{v}_{t,b}+\epsilon}\right)}$
$\leq$ $\sum_{\Phi(i)=b}\frac{\left|g_{t,i}\bar{g}_{t,i}\sqrt{1-\beta_{2}}\left(\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}}-\sqrt{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}\right)\right|}{\sqrt{v_{t,b}+\epsilon}\sqrt{\tilde{v}_{t,b}+\epsilon}}$
$\leq$ $\frac{1}{2}\sum_{\Phi(i)=b}\frac{\bar{g}_{t,i}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\frac{\left(\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}}-\sqrt{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}\right)^{2}}{\mathbb{E}[\left(\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}}-\sqrt{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}\right)^{2}|{\bm{x}}_{t-1}]}$
$+$ $\frac{1}{2}\sum_{\Phi(i)=b}\frac{(1-\beta_{2})g_{t,i}^{2}\mathbb{E}[\left(\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}}-\sqrt{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}\right)^{2}|{\bm{x}}_{t-1}]}{(v_{t,b}+\epsilon)\sqrt{\tilde{v}_{t,b}+\epsilon}}$
The first inequality is because $v_{t,b}+\epsilon\geq(1-\beta_{2})\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}$ and $\tilde{v}_{t,b}+\epsilon\geq(1-\beta_{2})\left(\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}\right)$. For the first term, it will be exactly $\frac{1}{2}\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}$ after taking expectation conditional on ${\bm{x}}_{t-1}$. For the second term, we have the following inequality
$\mathbb{E}\left[\left(\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}}-\sqrt{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}\right)^{2}\middle|{\bm{x}}_{t-1}\right]$
$=$ $\mathbb{E}\left[\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}+\sum_{\Phi(j)=b}g_{t,j}^{2}/d_{b}-2\sqrt{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{i}^{2}}\middle|{\bm{x}}_{t-1}\right]$
$\leq$ $2\left(\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}\right)-2\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}}\mathbb{E}\left[\sqrt{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}\middle|{\bm{x}}_{t-1}\right]$
$\leq$ $2\left(\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}\right)-2\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}}\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}}$
$=$ $2\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}}\left(\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}}-\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}}\right)$
$\leq$ $2\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}}\sigma_{b}.$
The first inequality comes from 3.4. The second inequality is because $\ell_{2}$ norm is a convex function. Then we know that
$\sum_{\Phi(i)=b}\frac{(1-\beta_{2})g_{t,i}^{2}\mathbb{E}\left[\left(\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}}-\sqrt{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}\right)^{2}|{\bm{x}}_{t-1}\right]}{(v_{t,b}+\epsilon)\sqrt{\tilde{v}_{t,b}+\epsilon}}$
$\leq$ $\sum_{\Phi(i)=b}\frac{(1-\beta_{2})g_{t,i}^{2}2\sqrt{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}}\sigma_{b}}{(v_{t,b}+\epsilon)\sqrt{\tilde{v}_{t,b}+\epsilon}}$
$\leq$ $2\sqrt{1-\beta_{2}}\sigma_{b}\sum_{\Phi(i)=b}\frac{g_{t,i}^{2}}{v_{t,b}+\epsilon}.$
Then back to Equation 10, we have that
$\mathbb{E}\sum_{t=1}^{T}\sum_{\Phi(i)=b}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,b}+\epsilon}}$ $=\mathbb{E}\sum_{t=1}^{T}\sum_{\Phi(i)=b}\frac{\bar{g}_{t,i}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}+\mathbb{E}\left[\sum_{t=1}^{T}\sum_{\Phi(i)=b}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,b}+\epsilon}}-\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\right]$
$\geq\mathbb{E}\sum_{t=1}^{T}\sum_{\Phi(i)=b}\frac{\bar{g}_{t,i}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}-\frac{1}{2}\mathbb{E}\sum_{t=1}^{T}\sum_{\Phi(i)=b}\frac{\bar{g}_{t,i}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}-\frac{1}{2}2\sqrt{1-\beta_{2}}\sigma_{b}\mathbb{E}\sum_{t=1}^{T}\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}}{v_{t,b}+\epsilon}$
$=\frac{1}{2}\mathbb{E}\sum_{t=1}^{T}\sum_{\Phi(i)=b}\frac{\bar{g}_{t,i}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}-\sqrt{1-\beta_{2}}\sigma_{b}\mathbb{E}\sum_{t=1}^{T}\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}}{v_{t,b}+\epsilon}$
For the second term, we can apply Lemma 3.13 and get that
$\sum_{t=1}^{T}\frac{\sum_{\Phi(i)=b}g_{t,i}^{2}/d_{b}}{v_{t,b}+\epsilon}\leq T+\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}.$
Combining these two terms, we can get that
$\mathbb{E}\sum_{t=1}^{T}\sum_{\Phi(i)=b}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,b}+\epsilon}}$ $\geq\frac{1}{2}\mathbb{E}\sum_{t=1}^{T}\sum_{\Phi(i)=b}\frac{\bar{g}_{t,i}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}-\sqrt{1-\beta_{2}}Td_{b}\sigma_{b}-\frac{d_{b}\sigma_{b}\beta_{2}}{\sqrt{1-\beta_{2}}}\mathbb{E}\left[\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}\right].$
Next we need Lemma 3.15 to deal with the denominator in the approximated first order term. The lemma is largely inspired by Lemma 6 in Li & Lin (2024), where we further generalize it to the case of block-wise $\mathtt{Adam}$.
See 3.15
Proof of Lemma 3.15.
For each $t\leq T$, we have that
$\mathbb{E}\left[\sqrt{\tilde{v}_{t,b}+\epsilon}\right]$
$=$ $\mathbb{E}\left[\sqrt{\beta_{2}v_{t-1,b}+(1-\beta_{2})(\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2})+\epsilon}\right]$
$=$ $\mathbb{E}\left[\frac{\beta_{2}v_{t-1,b}+(1-\beta_{2})\sigma_{b}^{2}+\epsilon}{\sqrt{\beta_{2}v_{t-1,b}+(1-\beta_{2})(\sum_{\Phi(i)=b}\bar{g}_{t,i}^{2}/d_{b}+\sigma_{b}^{2})+\epsilon}}\right]+(1-\beta_{2})\mathbb{E}\left[\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\right]$
$\leq$ $\mathbb{E}\left[\sqrt{\beta_{2}v_{t-1,b}+(1-\beta_{2})\sigma_{b}^{2}+\epsilon}\right]+(1-\beta_{2})\mathbb{E}\left[\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\right].$
And for each $s\leq t-1$, we have that
$\mathbb{E}\left[\sqrt{\beta_{2}^{s}v_{t-s,b}+(1-\beta_{2}^{s})\sigma_{b}^{2}+\epsilon}\right]$
$=$ $\mathbb{E}\left[\sqrt{\beta_{2}^{s+1}v_{t-s-1,b}+\beta_{2}^{s}(1-\beta_{2})\sum_{\Phi(i)=b}g_{t-s,i}^{2}/d_{b}+(1-\beta_{2}^{s})\sigma_{b}^{2}+\epsilon}\right]$
$=$ $\mathbb{E}\left[\mathbb{E}\left[\sqrt{\beta_{2}^{s+1}v_{t-s-1,b}+\beta_{2}^{s}(1-\beta_{2})\sum_{\Phi(i)=b}g_{t-s,i}^{2}/d_{b}+(1-\beta_{2}^{s})\sigma_{b}^{2}+\epsilon}\middle|{\bm{x}}_{t-s-1}\right]\right]$
$\leq$ $\mathbb{E}\left[\sqrt{\beta_{2}^{s+1}v_{t-s-1,b}+\beta_{2}^{s}(1-\beta_{2})\mathbb{E}\left[\sum_{\Phi(i)=b}g_{t-s,i}^{2}/d_{b}\middle|{\bm{x}}_{t-s-1}\right]+(1-\beta_{2}^{s})\sigma_{b}^{2}+\epsilon}\right]$
$\leq$ $\mathbb{E}\left[\sqrt{\beta_{2}^{s+1}v_{t-s-1,b}+\beta_{2}^{s}(1-\beta_{2})\sum_{\Phi(i)=b}\bar{g}_{t-s,i}^{2}/d_{b}+(1-\beta_{2}^{s+1})\sigma_{b}^{2}+\epsilon}\right]$
$=$ $\mathbb{E}\left[\frac{\beta_{2}^{s+1}v_{t-s-1,b}+(1-\beta_{2}^{s+1})\sigma_{b}^{2}+\epsilon}{\sqrt{\beta_{2}^{s+1}v_{t-s-1,b}+\beta_{2}^{s}(1-\beta_{2})\sum_{\Phi(i)=b}\bar{g}_{t-s,i}^{2}/d_{b}+(1-\beta_{2}^{s+1})\sigma_{b}^{2}+\epsilon}}\right]$
$+$ $\mathbb{E}\left[\frac{\beta_{2}^{s}(1-\beta_{2})\sum_{\Phi(i)=b}\bar{g}_{t-s,i}^{2}/d_{b}}{\sqrt{\beta_{2}^{s+1}v_{t-s-1,b}+\beta_{2}^{s}(1-\beta_{2})\sum_{\Phi(i)=b}\bar{g}_{t-s,i}^{2}/d_{b}+(1-\beta_{2}^{s+1})\sigma_{b}^{2}+\epsilon}}\right]$
$\leq$ $\mathbb{E}\left[\sqrt{\beta_{2}^{s+1}v_{t-s-1,b}+(1-\beta_{2}^{s+1})\sigma_{b}^{2}+\epsilon}\right]+\sqrt{\beta_{2}^{s}}(1-\beta_{2})\mathbb{E}\left[\frac{\sum_{\Phi(i)=b}\bar{g}_{t-s,i}^{2}/d_{b}}{\sqrt{\tilde{v}_{t-s,b}+\epsilon}}\right].$
By summing the above inequality over $s=1,\cdots,t-1$, we have that
$\mathbb{E}\left[\sqrt{\beta_{2}v_{t-1,b}+(1-\beta_{2})\sigma_{b}^{2}+\epsilon}\right]$
$\leq$ $\mathbb{E}\left[\sqrt{\beta_{2}^{t}v_{0,b}+(1-\beta_{2}^{t})\sigma_{b}^{2}+\epsilon}\right]+\sum_{s=1}^{t-1}\sqrt{\beta_{2}^{s}}(1-\beta_{2})\mathbb{E}\left[\frac{\sum_{\Phi(i)=b}\bar{g}_{t-s,i}^{2}/d_{b}}{\sqrt{\tilde{v}_{t-s,b}+\epsilon}}\right]$
$\leq$ $\sqrt{\beta_{2}^{t}v_{0,b}}+\sqrt{\sigma_{b}^{2}+\epsilon}+\sum_{s=1}^{t-1}\sqrt{\beta_{2}^{s}}(1-\beta_{2})\mathbb{E}\left[\frac{\sum_{\Phi(i)=b}\bar{g}_{t-s,i}^{2}/d_{b}}{\sqrt{\tilde{v}_{t-s,b}+\epsilon}}\right].$
and
$\mathbb{E}\left[\sqrt{\tilde{v}_{t,b}+\epsilon}\right]\leq\sqrt{\beta_{2}^{t}v_{0,b}}+\sqrt{\sigma_{b}^{2}+\epsilon}+\sum_{s=0}^{t-1}\sqrt{\beta_{2}^{s}}(1-\beta_{2})\mathbb{E}\left[\frac{\sum_{\Phi(i)=b}\bar{g}_{t-s,i}^{2}/d_{b}}{\sqrt{\tilde{v}_{t-s,b}+\epsilon}}\right].$
By summing the above inequality over $t=\frac{T}{2}+1,\cdots,T$, we have that
$\sum_{t=\frac{T}{2}+1}^{T}\left[\sqrt{\tilde{v}_{t,b}+\epsilon}\right]$ $\leq\sum_{t=\frac{T}{2}+1}^{T}\sqrt{\beta_{2}^{t}v_{0,b}}+\frac{T}{2}\sqrt{\sigma_{b}^{2}+\epsilon}+\sum_{t=\frac{T}{2}+1}^{T}\sum_{s=0}^{t-1}\sqrt{\beta_{2}^{s}}(1-\beta_{2})\mathbb{E}\left[\frac{\sum_{\Phi(i)=b}\bar{g}_{t-s,i}^{2}/d_{b}}{\sqrt{\tilde{v}_{t-s,b}+\epsilon}}\right]$
$\leq\frac{\beta_{2}^{\frac{T}{4}}}{1-\sqrt{\beta_{2}}}\sqrt{v_{0,b}}+\frac{T}{2}\sqrt{\sigma_{b}^{2}+\epsilon}+\frac{1-\beta_{2}}{1-\sqrt{\beta_{2}}}\sum_{t=1}^{T}\mathbb{E}\left[\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\right]$
$=\frac{\beta_{2}^{\frac{T}{4}}}{1-\sqrt{\beta_{2}}}\sqrt{v_{0,b}}+\frac{T}{2}\sqrt{\sigma_{b}^{2}+\epsilon}+(1+\sqrt{\beta_{2}})\sum_{t=1}^{T}\mathbb{E}\left[\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\right]$
$\leq\frac{2\beta_{2}^{\frac{T}{4}}}{1-\beta_{2}}\sqrt{v_{0,b}}+\frac{T}{2}\sigma_{b}+\frac{T}{2}\sqrt{\epsilon}+2\sum_{t=1}^{T}\mathbb{E}\left[\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\right].$
This following Lemma C.1 is to control the growth of $v_{T,b}$ so that the right hand side in Lemma 3.13 is indeed $\Theta\left(T+\frac{\log{T}}{1-\beta_{2}}\right)$ instead of $\Theta(T)$ when all the constants are $\text{poly}(T)$.
Lemma C.1.
For any $T$, it holds that
$\ln\frac{\mathbb{E}\max_{b\in[B]}v_{T,b}+\epsilon}{v_{0}+\epsilon}$ $\leq 2\ln\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}+\max_{b\in[B]}H_{b}^{2}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}$
Proof of Lemma C.1.
From the definition of $v_{t,i}$ and 3.4, we have that
$\mathbb{E}\max_{b\in[B]}v_{t,b}$
$=$ $\mathbb{E}\max_{b\in[B]}\left[\beta_{2}^{t}v_{0,b}+(1-\beta_{2})\sum_{s=1}^{t}\beta_{2}^{t-s}\left\|{\bm{g}}_{s,(b)}\right\|_{2}^{2}/d_{b}\right]$
$\leq$ $\beta_{2}^{t}\left\|{\bm{v}}_{0}\right\|_{\infty}+(1-\beta_{2})\mathbb{E}\max_{b\in[B]}\sum_{s=1}^{t}\beta_{2}^{t-s}\left\|{\bm{g}}_{s,(b)}\right\|_{2}^{2}/d_{b}$
$=$ $\beta_{2}^{t}\left\|{\bm{v}}_{0}\right\|_{\infty}+(1-\beta_{2})\mathbb{E}\max_{b\in[B]}\sum_{s=1}^{t}\beta_{2}^{t-s}\left\|\mathbb{E}[{\bm{g}}_{s,(b)}|{\bm{x}}_{s-1}]+{\bm{g}}_{s,(b)}-\mathbb{E}[{\bm{g}}_{s,(b)}|{\bm{x}}_{s-1}]\right\|_{2}^{2}/d_{b}$
$\leq$ $\beta_{2}^{t}\left\|{\bm{v}}_{0}\right\|_{\infty}+(1-\beta_{2})\mathbb{E}\max_{b\in[B]}\sum_{s=1}^{t}\beta_{2}^{t-s}\left[2\left\|\mathbb{E}[{\bm{g}}_{s,(b)}|{\bm{x}}_{s-1}]\right\|_{2}^{2}+2\left\|{\bm{g}}_{s,(b)}-\mathbb{E}[{\bm{g}}_{s,(b)}|{\bm{x}}_{s-1}]\right\|_{2}^{2}\right]/d_{b}$
$\leq$ $\beta_{2}^{t}\left\|{\bm{v}}_{0}\right\|_{\infty}+2(1-\beta_{2})\mathbb{E}\sum_{b=1}^{B}\sum_{s=1}^{t}\beta_{2}^{t-s}\left\|{\bm{g}}_{s,(b)}-\mathbb{E}[{\bm{g}}_{s,(b)}|{\bm{x}}_{s-1}]\right\|_{2}^{2}/d_{b}+2(1-\beta_{2})\mathbb{E}\max_{b\in[B]}\sum_{s=1}^{t}\beta_{2}^{t-s}\left\|\nabla_{(b)}L({\bm{x}}_{s-1})\right\|_{2}^{2}/d_{b}$
$\leq$ $\beta_{2}^{t}\left\|{\bm{v}}_{0}\right\|_{\infty}+2(1-\beta_{2}^{t})\sum_{b=1}^{B}\sigma_{b}^{2}+2(1-\beta_{2})\mathbb{E}\max_{b\in[B]}\sum_{s=1}^{t}\beta_{2}^{t-s}\left[2\left\|\nabla_{(b)}L({\bm{x}}_{0})\right\|_{2}^{2}+2\left\|\nabla_{(b)}L({\bm{x}}_{s-1})-\nabla_{(b)}L({\bm{x}}_{0})\right\|_{2}^{2}\right]/d_{b}$
$\leq$ $\beta_{2}^{t}\left\|{\bm{v}}_{0}\right\|_{\infty}+2(1-\beta_{2}^{t})\sum_{b=1}^{B}\sigma_{b}^{2}+4(1-\beta_{2}^{t})\max_{b\in[B]}\left\|\nabla_{(b)}L({\bm{x}}_{0})\right\|_{2}^{2}/d_{b}$
$+$ $4(1-\beta_{2})\mathbb{E}\max_{b\in[B]}\sum_{s=1}^{t}\beta_{2}^{t-s}\left\|\nabla_{(b)}L({\bm{x}}_{s-1})-\nabla_{(b)}L({\bm{x}}_{0})\right\|_{2}^{2}/d_{b}$
$\leq$ $\beta_{2}^{t}\left\|{\bm{v}}_{0}\right\|_{\infty}+2(1-\beta_{2}^{t})\sum_{b=1}^{B}\sigma_{b}^{2}+4(1-\beta_{2}^{t})\max_{b\in[B]}\left\|\nabla_{(b)}L({\bm{x}}_{0})\right\|_{2}^{2}/d_{b}$
$+$ $4(1-\beta_{2})\mathbb{E}\max_{b\in[B]}\sum_{s=1}^{t}\beta_{2}^{t-s}\frac{H_{b}^{2}}{d_{b}^{2}}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{s-1,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}$
$\leq$ $\beta_{2}^{t}\left\|{\bm{v}}_{0}\right\|_{\infty}+2\sum_{b=1}^{B}\sigma_{b}^{2}+4\max_{b\in[B]}\left\|\nabla_{(b)}L({\bm{x}}_{0})\right\|_{2}^{2}/d_{b}+4(1-\beta_{2})(\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}})\mathbb{E}\sum_{s=1}^{t}\beta_{2}^{t-s}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{s-1,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}.$
We define $C=v_{0}+2\sum_{b=1}^{B}\sigma_{b}^{2}+4\max_{b\in[B]}\left\|\nabla_{(b)}L({\bm{x}}_{0})\right\|_{2}^{2}/d_{b}$ for simplicity. From Lemma 3.13 and Cauchy inequality, we know that
$\frac{1}{d_{b^{\prime}}}\left\|{\bm{x}}_{t,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}$ $=\frac{\eta^{2}}{d_{b^{\prime}}}\sum_{\Phi(j)=b^{\prime}}\left|\sum_{s=1}^{t}\frac{g_{s,j}}{\sqrt{v_{s,b^{\prime}}+\epsilon}}\right|^{2}$
$\leq\frac{\eta^{2}}{d_{b^{\prime}}}\sum_{\Phi(j)=b^{\prime}}t\sum_{s=1}^{t}\frac{g_{s,j}^{2}}{v_{s,b^{\prime}}+\epsilon}$
$=\eta^{2}t\sum_{s=1}^{t}\frac{\sum_{\Phi(j)=b^{\prime}}g_{s,j}^{2}/d_{b^{\prime}}}{v_{s,b^{\prime}}+\epsilon}$
$\leq\eta^{2}t\left(t+\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{v_{t,b^{\prime}}+\epsilon}{v_{0,b^{\prime}}+\epsilon}\right)$
$\leq\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\max_{b\in[B]}v_{t,b}+\epsilon}{v_{0}+\epsilon}$
The right hand side is independent of specific block $b$. So we can get that
$\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{t,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}\leq\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\max_{b\in[B]}v_{t,b}+\epsilon}{v_{0}+\epsilon}$
and
$\mathbb{E}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{t,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}$
$\leq$ $\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\mathbb{E}\ln\frac{\max_{b\in[B]}v_{t,b}+\epsilon}{v_{0}+\epsilon}$
$\leq$ $\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\mathbb{E}\max_{b\in[B]}v_{t,b}+\epsilon}{v_{0}+\epsilon}$
$\leq$ $\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\epsilon+C+4(1-\beta_{2})\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\mathbb{E}\sum_{s=1}^{t}\beta_{2}^{t-s}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{s-1,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}}{v_{0}+\epsilon}.$
Define $G=\max_{1\leq t\leq T}\mathbb{E}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{t,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}$. There exists $t\leq T$ such that
$G$ $=\mathbb{E}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{t,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}$
$\leq\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\epsilon+C+4(1-\beta_{2})\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\mathbb{E}\sum_{s=1}^{t}\beta_{2}^{t-s}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{s-1,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}}{v_{0}+\epsilon}$
$\leq\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\epsilon+C+4(1-\beta_{2})\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\sum_{s=1}^{t}\beta_{2}^{t-s}G}{v_{0}+\epsilon}$
$\leq\eta^{2}T^{2}+\eta^{2}T\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\epsilon+C+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}G}{v_{0}+\epsilon}.$
We further define $G^{\prime}=\epsilon+C+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}G$ and get that
$G^{\prime}$ $\leq\epsilon+C+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T^{2}+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{G^{\prime}}{v_{0}+\epsilon}}$
$\leq\epsilon+C+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T^{2}$
$\phantom{\leq\epsilon}+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{1-\beta_{2}}\left(\ln{\frac{G^{\prime}(1-\beta_{2})}{4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\beta_{2}}}+\ln{4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}}\right)$
$\leq\epsilon+C+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T^{2}+\frac{G^{\prime}}{2}+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{1-\beta_{2}}\ln{4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}}$
$\leq\epsilon+C+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T^{2}+\frac{G^{\prime}}{2}+(v_{0}+\epsilon)\left(4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}\right)^{2}.$
The last two inequalites come from $\ln{x}\leq\frac{x}{2}$ and $x\ln(x)\leq x^{2}$. Then we can get that
$\mathbb{E}\max_{b\in[B]}v_{T,b}+\epsilon$ $\leq\epsilon+C+4(1-\beta_{2})\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\sum_{s=1}^{t}\beta_{2}^{t-s}G$
$\leq\epsilon+C+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}G=G^{\prime}$
$\leq 2\epsilon+2C+8\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T^{2}+32(v_{0}+\epsilon)\left(\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}\right)^{2}$
and
$\ln{\frac{\mathbb{E}\max_{b\in[B]}v_{T,b}+\epsilon}{v_{0}+\epsilon}}$
$\leq$ $\ln{\frac{2\epsilon+2C+8\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T^{2}+32(v_{0}+\epsilon)\left(\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}\right)^{2}}{v_{0}+\epsilon}}$
$\leq$ $\ln{\left[2\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+2\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}}{v_{0}+\epsilon}+\frac{2\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T^{2}}{v_{0}+\epsilon}+\frac{4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}\right)^{2}\right]}$
$\leq$ $2\ln\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}+\max_{b\in[B]}H_{b}^{2}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}$
Finally, we give the proof for Theorem 3.11. When $\Phi(i)=i$, i.e., each parameter forms a single block, it becomes the proof for Theorem 3.5. See 3.11
Proof of Theorem 3.11.
In a single step, we can apply Lemma 3.12 and have that
$L({\bm{x}}_{t})-L({\bm{x}}_{t-1})$ $\leq\nabla L({\bm{x}}_{t-1})^{\top}({\bm{x}}_{t}-{\bm{x}}_{t-1})+\frac{1}{2}\sum_{b=1}^{B}\frac{H_{b}}{d_{b}}\sum_{\Phi(i)=b}\left(x_{t,i}-x_{t-1,i}\right)^{2}$
$=-\eta\sum_{i=1}^{d}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,\Phi(i)}+\epsilon}}+\frac{1}{2}\eta^{2}\sum_{b=1}^{B}\frac{H_{b}}{d_{b}}\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}}{v_{t,b}+\epsilon}.$
If we sum over $t$ from $1$ to $T$ and take expectation, we can get
$\mathbb{E}\left[L({\bm{x}}_{T})-L({\bm{x}}_{0})\right]$ $\leq-\mathbb{E}\left[\eta\sum_{i=1}^{d}\sum_{t=1}^{T}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,\Phi(i)}+\epsilon}}\right]+\frac{1}{2}\eta^{2}\mathbb{E}\left[\sum_{b=1}^{B}\frac{H_{b}}{d_{b}}\sum_{t=1}^{T}\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}}{v_{t,b}+\epsilon}\right]$
$\leq-\mathbb{E}\left[\eta\sum_{i=1}^{d}\sum_{t=1}^{T}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,\Phi(i)}+\epsilon}}\right]+\frac{1}{2}\eta^{2}\mathbb{E}\left[\sum_{b=1}^{B}H_{b}\left(T+\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}\right)\right].$
The second inequality comes from applying Lemma 3.13. By Lemma 3.14, we have that
$\frac{1}{T}\mathbb{E}\left[\sum_{i=1}^{d}\sum_{t=1}^{T}\frac{\bar{g}_{t,i}^{2}}{\sqrt{\tilde{v}_{t,\Phi(i)}+\epsilon}}\right]$ $\leq\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\frac{\eta}{T}\mathbb{E}\left[\sum_{b=1}^{B}H_{b}\left(T+\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}\right)\right]$
$+\frac{1}{T}\sum_{b=1}^{B}d_{b}\sigma_{b}\sqrt{1-\beta_{2}}\left(T+\frac{\beta_{2}}{1-\beta_{2}}\mathbb{E}\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}\right)$
$\leq\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\eta\sum_{b=1}^{B}H_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}$
$+\frac{\beta_{2}}{T(1-\beta_{2})}\left(\eta\sum_{b=1}^{B}H_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}\sigma_{b}\right)\max_{b\in[B]}\mathbb{E}\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}$
$\leq\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\eta\sum_{b=1}^{B}H_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}$
$+\frac{\beta_{2}}{T(1-\beta_{2})}\left(\eta\sum_{b=1}^{B}H_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}\right)\ln{\frac{\mathbb{E}\max_{b\in[B]}v_{T,b}+\epsilon}{v_{0}+\epsilon}}$
From Lemma C.1, we can define
$E$ $=\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\eta\sum_{b=1}^{B}H_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}+\frac{\beta_{2}}{T(1-\beta_{2})}\left(\eta\sum_{b=1}^{B}H_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}\right)F,$
with
$F=2\ln\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}+\max_{b\in[B]}H_{b}^{2}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}.$
Then it holds that
$\frac{1}{T}\mathbb{E}\left[\sum_{i=1}^{d}\sum_{t=1}^{T}\frac{\bar{g}_{t,i}^{2}}{\sqrt{\tilde{v}_{t,\Phi(i)}+\epsilon}}\right]\leq E.$
By Lemma 3.15 and Cauchy inequality, we have that
$\frac{2}{T}\mathbb{E}\sum_{t=\frac{T}{2}+1}^{T}\sum_{b=1}^{B}\sqrt{d_{b}}\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}$ $\leq\left(\frac{2}{T}\mathbb{E}\sum_{t=\frac{T}{2}+1}^{T}\sum_{b=1}^{B}\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\right)^{\frac{1}{2}}\left(\frac{2}{T}\mathbb{E}\sum_{t=\frac{T}{2}+1}^{T}\sum_{b=1}^{B}d_{b}\sqrt{\tilde{v}_{t,b}+\epsilon}\right)^{\frac{1}{2}}$
$\leq\left(\frac{2}{T}\mathbb{E}\sum_{t=\frac{T}{2}+1}^{T}\sum_{b=1}^{B}\sum_{\Phi(i)=b}\frac{\bar{g}_{t,i}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}\right)^{\frac{1}{2}}\left(\frac{2}{T}\mathbb{E}\sum_{t=\frac{T}{2}+1}^{T}\sum_{b=1}^{B}d_{b}\sqrt{\tilde{v}_{t,b}+\epsilon}\right)^{\frac{1}{2}}$
$\leq\sqrt{2E}\left(4E+\frac{4\beta_{2}^{\frac{T}{4}}}{T(1-\beta_{2})}d\sqrt{v_{0}}+\sum_{b=1}^{B}d_{b}\sigma_{b}+d\sqrt{\epsilon}\right)^{\frac{1}{2}}$
$\leq 2\sqrt{2}E+\sqrt{2}\sqrt{E}\sqrt{\frac{4\beta_{2}^{\frac{T}{4}}}{T(1-\beta_{2})}d\sqrt{v_{0}}+\sum_{b=1}^{B}d_{b}\sigma_{b}+d\sqrt{\epsilon}}.$
This completes the proof. ∎
Appendix D Experiment Details
D.1 $\mathtt{Adam}$ on a rotated loss
A key difficulty in implementing $\mathtt{rotated\ Adam}$ arises from applying an orthogonal rotation on the parameters before calculating the loss. It is computationally infeasible to apply a 125M $\times$ 125M orthogonal matrix on the 125M-sized parameter vector. To avoid such computation, we design a new orthogonal transformer to rotate the parameters of the network. In what follows, we elaborate on this rotation.
$\mathtt{RandPerm}$.
Given a vector $v$ of size $d$, we can orthogonally rotate it by repeatedly applying these consecutive operations: 1. Permute the entries of the vector according to a randomly chosen permutation $\pi\in\mathbb{S}_{d}$. 2. Reshape the permuted vector into a 3D tensor of size $[s_{1},s_{2},s_{3}]$, apply a fixed orthogonal rotation of size $s\times s$ on each side of the tensor and then reshape it back to a vector of size $d$.
[…]
D.2 Computation of matrix norms
It is computationally infeasible to get the full Hessian matrix and directly compute norms of it. Instead we leverage Hessian vector product function in Jax to probe the Hessian matrix. We use Lanczos algorithm to estimate spectral norm. The estimation of $(1,1)$-norm relies on the properties of Cauchy distribution. Given $a=[a_{1},\cdots,a_{n}]$ and i.i.d. standard Cauchy variables $X_{1},\cdots,X_{n}$, $\sum_{i=1}^{n}a_{i}X_{i}$ is Cauchy distributed with mean 00 and scale $\sum_{i=1}^{n}\left|a_{i}\right|$. For a single value Cauchy distribution with mean 00 and scaling $\gamma$, an estimator for $\gamma$ is the median of the absolute values of all the samples.
Therefore, we propose Algorithm 4 to estimate the sum of absolute values for each row and sum over all the rows to get $(1,1)$-norm of Hessian matrix. We choose $n=50$ for all the measurement experiments. We also prove a concentration inequality in Theorem D.1.
Algorithm 4 Estimation of $(1,1)$-Norm of Hessian, $\nabla^{2}L(\theta)$
1:Number of Cauchy vectors $n$, parameter $\theta\in\mathbb{R}^{d}$, loss $L$
2:for $i=1\to n$ :
[…]
4:   $\mathbf{H}_{:,i}\leftarrow\nabla^{2}L(\theta)\cdot v^{(i)}$ (Using hessian-vector product)
5:return $\sum_{j=1}^{d}\mathrm{median}(\mathrm{abs}(\mathbf{H}_{j,:}))$
Theorem D.1.
For the estimate in Algorithm 4, it holds that
$P\left(\left|\sum_{j=1}^{d}\textup{median}(\left|\mathbf{H}_{j,:}\right|)-\left\|\nabla^{2}L(\theta)\right\|_{1,1}\right|\geq\epsilon\left\|\nabla^{2}L(\theta)\right\|_{1,1}\right)\leq 2d\cdot\exp{\left(-\frac{8n}{25\pi^{2}}\epsilon^{2}\right)}$
for every $\epsilon\in(0,1)$.
Proof of Theorem D.1.
[…]
$P(\left|M_{n}-1\right|\geq\epsilon)\leq 2\exp{\left(-\frac{8n}{25\pi^{2}}\epsilon^{2}\right)}.$ (11)
Given $X\sim\textrm{Cauchy}[0,1]$, the cumulative distribution function of $\left|X\right|$ is $F(x)=\frac{2}{\pi}\textrm{arctan}(x)$ and the median of $\left|X\right|$ is $1$ because $F(1)=0.5$. For a fixed $\epsilon\in(0,1)$, define $p_{1}=\frac{2}{\pi}\textrm{arctan}(1-\epsilon)$ to be the probability that $\left|X\right|$ is smaller than $1-\epsilon$ and $S=\sum_{i=1}^{n}\mathbf{1}_{\left|X_{i}\right|\leq 1-\epsilon}$. Since $\mathbf{1}_{\left|X_{i}\right|\leq 1-\epsilon}$ follows i.i.d. Bernoulli distribution with $p_{1}$, $S\sim\textrm{Bin}(n,p_{1})$.
$M_{n}\leq 1-\epsilon$ if and only if at least $\frac{n+1}{2}$ $X_{i}$’s are smaller than $1-\epsilon$. And we can apply Hoeffding’s inequality on $S$ and get that
$P(M_{n}\leq 1-\epsilon)$ $=P(S\geq\frac{n+1}{2})\leq P(S\geq\frac{n}{2})$
$=P(S-np_{1}\geq\frac{n}{2}-np_{1})$
$\leq\exp{\left(-\frac{2(\frac{n}{2}-np_{1})^{2}}{n}\right)}$
$=\exp{\left(-2n\left(\frac{1}{2}-\frac{2}{\pi}\textrm{arctan}(1-\epsilon)\right)^{2}\right)}.$
With similar derivation, we can also get that
$P(M_{n}\geq 1+\epsilon)\leq\exp{\left(-2n\left(\frac{1}{2}-\frac{2}{\pi}\textrm{arctan}(1+\epsilon)\right)^{2}\right)}.$
When $\epsilon\in[-1,1]$, we have that
$\left(\frac{1}{2}-\frac{2}{\pi}\textrm{arctan}(1+\epsilon)\right)^{2}=\frac{4}{\pi^{2}}\left(\textrm{arctan}(1)-\textrm{arctan}(1+\epsilon)\right)^{2}\geq\frac{4}{25\pi^{2}}\epsilon^{2}.$
because $\textrm{arctan}^{\prime}(x)=\frac{1}{1+x^{2}}$ is always greater than $\frac{1}{5}$ for $x\in[0,2]$. Then we can have Equation 11.
[…]
$P\left(\left|\textrm{median}(\left|H_{j,1}\right|,\cdots,\left|H_{j,n}\right|)-\sum_{k=1}^{d}\left|\nabla^{2}L(\theta)_{j,k}\right|\right|\geq\epsilon\sum_{k=1}^{d}\left|\nabla^{2}L(\theta)_{j,k}\right|\right)\leq 2\exp{\left(-\frac{8n}{25\pi^{2}}\epsilon^{2}\right)}.$
And we can sum the tail probability over all the rows $j$ and get that
$P\left(\left|\sum_{j=1}^{d}\textrm{median}(\left|\mathbf{H}_{j,:}\right|)-\left\|\nabla^{2}L(\theta)\right\|_{1,1}\right|\geq\epsilon\left\|\nabla^{2}L(\theta)\right\|_{1,1}\right)\leq 2d\exp{\left(-\frac{8n}{25\pi^{2}}\epsilon^{2}\right)}.$